### v1

# Vehicle Rebalancing Under Adherence Uncertainty

Abstract
Ride hailing systems suffer from spatial-temporal supply demand imbalance due to drivers operating in independent, and uncoordinated manner. Several fleet rebalancing models have been proposed that can provide repositioning recommendations to idle standing drivers with the objective of maximizing service rate or minimizing customer waiting time. Existing models assume complete adherence by the drivers which leads to limited practical implementation of these models. A novel taxi rebalancing model and a procedure to compute recommendations for repositioning of taxis are proposed which accounts for uncertainties in adherence arising from taxi driver’s preferences and the evolving confidence of the driver in the system due to outcomes of repositioning recommendations. Extensive simulations using NYC taxi dataset showed that the proposed model can lead to 6.3% higher allocation rates, 11.8% higher driver profits, 4.5% higher demand fulfillment, and 14% higher confidence of drivers in the rebalancing system, as compared to a state-of-the-art rebalancing model that is agnostic to the taxi driver preferences and confidence in the system.
Index Terms:
vehicle rebalancing, human factors in cyber-physical systems.
I Introduction
Recent urbanization trends have led to increased travel and its associated externalities, such as escalating congestion levels worldwide. In this context, reliance on private cars for personal mobility is becoming increasingly impractical and unsustainable. Ride-hailing services have emerged as an alternative, providing Mobility-on-Demand (MoD) services, but they raise concerns related to the exploitation of public resources, equity, profitability, and, importantly, scalability. Specifically, the travel demand for such services is spatio-temporally asymmetrically distributed (e.g., commuting toward downtown in the morning), making overall operations imbalanced and extremely sensitive to disturbances [1]. Fleet rebalancing has emerged as the solution to address these challenges. Two main classes of approaches have been proposed in the literature [1]. First, researchers have studied mathematical optimization models [2, 3, 4, 5, 6, 7, 8, 9], where upcoming demand is assumed to be known either exactly or approximated using demand prediction models. The optimization model then calculates rebalancing policies based on predicted demands. Second, researchers have employed reinforcement learning techniques to learn taxi rebalancing policies directly from data [10, 11, 12, 13]. Furthermore, from the decision-making perspective, models can be categorized into two classes, i) aggregate models [2, 3, 4, 5], and ii) agent-level models  [6, 7, 8, 9]. The former generate decisions at the fleet-level, determining how many taxis should be moved from one region to another. However, such models do not consider the state of individual taxi drivers, making it challenging to incorporate driver-level objectives such as fairness and preferences. In contrast, agent-level models possess the desired granularity, and offer greater flexibility and control over individual taxis, allowing for tailored movements based on specific criteria or constraints. However, the enhanced flexibility comes at a computational cost, due to the many variables involved in tracking each vehicle. One of the seminal works in the domain of the taxi rebalancing problem proposes an aggregate-level model for a station based taxi fleet management [2]. The objective is to find the rate at which idle taxis should be dispatched from one station to the other to ensure that the queue length of customers waiting at the station remains bounded. Based on the formulation proposed by the authors, the idle-standing taxis are distributed among the stations proportionally to the number of awaiting customers at each station. The excess taxis (if any) are evenly distributed among the stations at a minimum traveling cost. The primary drawback of this work is that it doesn’t consider the upcoming demands while generating the rebalancing decision, and the repositioning decisions are made solely based on the number of customers waiting in the queue. Subsequent work [3] introduced a more complex model, by accounting for both awaiting customers and expected future customer requests in the decision-making process. Taxis are assigned to stations to guarantee that each station has at least as many taxis as there are awaiting customers. Furthermore, following the allocation for awaiting customers, any remaining taxis are distributed proportionally based on demand forecasts at each station. A major shortcoming of [3] is that a simplistic travel demand forecasting assuming demand arrival as a Poisson process. Refer to [14] for a detailed analysis on travel demand forecasting. Secondly, [3] does not account for the uncertainties in the demand forecasts. This is critical, because recommendations might be overly optimistic and cause drivers to head to locations with insufficient demand. This may cause distrust among drivers and hence lead to poor adherence to system recommendations in the future. Authors in [4] further enhanced the aggregate-level taxi rebalancing solution by formulating a multi-period stochastic rebalancing problem considering uncertainties in demand forecasts. The objective of this framework is to ensure that the supply-demand ratio at each station is the same as the overall supply-demand ratio at the network level while minimizing the traveling cost. Formulating the taxi rebalancing problem as a multi-period problem has the advantage of accounting for the downstream effects of decisions made in the planning horizon. As pointed out by the authors, however, this may not always be beneficial due to the underlying uncertainties in the forecasts for mobility patterns. Authors in [5] discuss an upper-confidence-bound based approach to determine optimal parameters, such as the rebalancing frequency, and the length of the planning horizon, to ensure the effective implementation of the aggregate-level models sequentially.
All the models discussed so far are aggregate models and are unable to generate taxi-level decisions. [6] is the agent-level version of the aggregate model introduced by authors in [4], and enables decision-making at the individual taxi level. A notable limitation of this work is its failure to account for uncertainties in travel time and their potential impact on taxi availability at the destination. The authors in [7] greatly improved on this aspect and accounted for the impact of the travel time on taxi availability at the destination station. The availability of a taxi at the destination was represented as a fraction of the planning horizon during which the taxi is expected to be present at the destination, derermined using pre-calculated inter-station travel times. Although [6, 7] present agent-level models, they do not incorporate any driver-centric objectives, rather the objective still remains to balance network-level supply and demand. In our previous work [8], we improved on this aspect by incorporating fairness in providing rebalancing recommendations and [9] and [12] discussed the impact of the charging-related constraints on the taxi rebalancing problem. Apart from vehicle rebalancing. dynamic pricing has also been proposed as a method to mitigate the supply-demand imbalances [15] [16].
Interestingly, research on the impact of human factors on recommender systems is limited. In particular, little attention is given to the analysis of confidence level of taxi drivers and the impact of the rebalancing strategies suggested by the recommender systems on the adherence in the future. Existing taxi rebalancing models are agnostic to taxi driver’s confidence levels as well as their repositioning preferences. Due to this, the rebalancing recommendations given by the model may not be effective in practice because the driver may discard the recommendation if they are not confident about the recommender system or if the recommendation is not aligned with their preference. Driver’s confidence levels are impacted by the outcome of the repositioning recommendation: if the recommendation leads to a higher reward, the confidence level increases and vice-versa. Additionally, at any particular time instant, the confidence level impacts the likelihood of the driver accepting the repositioning recommendation from the system. The dynamics of an agent’s confidence while interacting with a recommender system has been discussed in literature. [17] proposed a Thompson Sampling-based decision-making and confidence update mechanism. This approach has been successfully applied in improving the exploration-exploitation trade-off in recommendation systems [18]. However, the implementation of such methods in taxi recommender systems remains limited. Prior works [19, 20, 21] have also shown that taxi drivers display a preference in terms of their passenger-finding strategies. [21] discussed the driver cruising strategies between the trips and highlights that taxi drivers tend to make reasonable choices between rebalancing and parking, heading to high-demand locations based on the time of day. Authors in [19] used a multinomial logit model to learn the probability distribution over the driver’s choice to head to one of the neighboring zones. Furthermore, authors in [20] presented a model that explores the behavior of taxi drivers when searching for passengers under uncertain conditions and three distinct strategies were studied: random search, maximum anticipated pick-up probability search, and maximum anticipated revenue search. A work that closely relates to our work is by  [22] which discusses a survey based driver preference aware repositioning recommendation model. A notable difference with this work is that it doesn’t consider the impact of the outcome of the repositioning recommendation on the driver’s probability of accepting the recommendation in the future and assumes a static acceptance probability. In our work, we model the driver’s confidence as a dynamic process, which is impacted by repeated interaction with the recommender system.
In this paper, we present a new detailed model for the behavior of taxi drivers, and tools to compute recommendations for such drivers, so that they can reposition their vehicles to increase their returns. Taxi drivers are modeled as agents characterized by a repositioning preference as well as a confidence level. We leverage the dataset reported in [23] to extract the taxi demands as well as driver preferences. A novel taxi rebalancing model has been proposed which provides repositioning recommendations by predicting the upcoming taxi demands, travel times, driver position, repositioning preference, and confidence in the recommender system. Repositioning recommendations are made sequentially and the agent’s confidence level is updated depending on the reward received after reaching the destination. To learn the taxi driver preference, we first identify the features that can describe the repositioning decision of a driver (e.g., the distance of the driver to the recommended destination, time of the day, day of the week, the expected number of requests at the destination, expected revenue from the ride originating at the destination, etc.). Next, we learn the probability of the driver moving to the destination using a logistic model and identify the top-k most likely destinations. When a driver is standing idle after dropping off a passenger, their next location is a function of the repositioning recommendation, their confidence level, and their preference. The confidence level of a taxi driver is modeled using the Thompson sampling-based Beta-Bernoulli Bandits, inspired by [17]. The performance of the confidence- and preference-aware repositioning model is compared with a state-of-the-art the agnostic repositioning model [7].
The key contributions of this work are detailed as follows:
  1. 1.
    We propose to learn the taxi driver preferences using logistic regression and model the taxi driver confidence as a dynamic process using Thompson sampling-based Beta-Bernoulli Bandits. To the best of our knowledge this is the first work to model the taxi drivers’ confidence in the recommender system and its impact on rebalancing recommendations.
  2. 2.
    We formulate and solve a novel taxi rebalancing optimization model that predicts the upcoming demands, travel times, driver positions, driver rebalancing preferences, and driver confidence, and uses that as an input to generate taxi rebalancing decisions to maximize profit.
  3. 3.
    We analyze the proposed strategy in a simulated taxi network derived from real taxi data from a state-of-the-art New York City dataset, and illustrate its effectiveness in comparison to preference and confidence-agnostic state-of-the-art taxi rebalancing model.
The rest of this manuscript is organized as follows. In Section II a rigorous mathematical formulation of the proposed preference- and confidence-aware rebalancing model is introduced, followed by the detailed presentation of the proposed algorithm for the rebalancing recommender system in Section III. We present detailed simulation-based case studies in Section IV, and draw conclusions in Section V.
II Mathematical Formulation
[…]
Figure 1: This figure illustrates the sequential decision-making process of a taxi driver $c\in\mathcal{C}$, showing the driver’s decision to accept or reject a system recommendation, the subsequent movement to a region, and the allocation outcome.
In this section, we present a detailed mathematical formulation of the proposed taxi drivers’ state aware repositioning recommender system. The system consists of a set of regions denoted as $\mathcal{R}$, and a set of idle taxi drivers denoted as $\mathcal{C}$. Drivers transport the passengers from one region to the other. Idle standing drivers receive a repositioning recommendation which they can either accept or reject. The driver moves to the recommended region if he accepts the system recommendation, or moves to his preferred region if he rejects the system recommendation. Upon reaching, the driver either gets allocated to transport a passenger, or remains idle.
This sequential decision-making process for driver $c\in\mathcal{C}$ is modelled as a random experiment with a probability space $(\Omega,\mathcal{F},\mathbb{P}_{c})$ and is shown in Fig. 1. The sample space of the random experiment is $\Omega=\{0,1\}\times\mathcal{R}\times\{0,1\}$, where an outcome is a tuple $\omega=(\omega_{1},\omega_{2},\omega_{3})\in\Omega$. Here, $\omega_{1}=0$ indicates that the driver rejected the system recommendation, while $\omega_{1}=1$ indicates that the driver accepted the system recommendation. Additionally, $\omega_{2}=j$ implies that the driver moved to region $j\in\mathcal{R}$. Finally, $\omega_{3}=1$ implies that the driver was allocated, whereas $\omega_{3}=0$ indicates that the driver remained idle. The $\sigma$-algebra, of the sample space is denoted as $\mathcal{F}$. For brevity, the following notation is used: $\{\omega_{1}=a\}=\{(\omega_{1},\omega_{2},\omega_{3})\in\Omega\mid\omega_{1}=a\}$. Similarly, this notation can be extended to other combinations of $\omega_{1}$, $\omega_{2}$, and $\omega_{3}$ to represent events involving multiple conditions. The probability measure is defined such that $\mathbb{P}_{c}(\{\omega_{1}=1\})$ represents the probability of the driver accepting the system recommendation, while $\mathbb{P}_{c}(\{\omega_{1}=0\})$ represents the probability of the driver rejecting the system recommendation. Furthermore, $\mathbb{P}_{c}(\{\omega_{2}=j\}\mid\{\omega_{1}=0\})$ denotes the probability of the driver moving to destination $j\in\mathcal{R}$ given that the driver has rejected the system recommendation, and $\mathbb{P}_{c}(\{\omega_{2}=j\}\mid\{\omega_{1}=1\})$ denotes the probability of the driver moving to destination $j\in\mathcal{R}$ given that the driver has accepted the system recommendation. Finally, the probability of the driver getting allocated is defined such that if the driver rejected the system recommendation and moved to region $j\in\mathcal{R}$, the probability that the driver remains idle is $\mathbb{P}_{c}(\{\omega_{3}=0\}\mid\{\omega_{1}=0\},\{\omega_{2}=j\})=1-\theta_{p}$, and the probability that the driver gets allocated is $\mathbb{P}_{c}(\{\omega_{3}=1\}\mid\{\omega_{1}=0\},\{\omega_{2}=j\})=\theta_{p}$. Note that the value of $\theta_{p}$ is unique to the driver $c\in\mathcal{C}$. Similarly, if the driver accepted the system recommendation and moved to region $j\in\mathcal{R}$, the probability that the driver remains idle is $\mathbb{P}_{c}(\{\omega_{3}=0\}\mid\{\omega_{1}=1\},\{\omega_{2}=j\})=1-\theta_{r}$, and the probability that the driver gets allocated is $\mathbb{P}_{c}(\{\omega_{3}=1\}\mid\{\omega_{1}=1\},\{\omega_{2}=j\})=\theta_{r}$.
The state of each taxi driver is defined as tuple containing i) probability of the driver going to region $j\in\mathcal{R}$ by following his own preference, and ii) probability of the driver accepting the reposition recommendation from the system.
$\left(\mathbb{P}_{c}({\omega_{2}=j|\omega_{1}=0}),\mathbb{P}_{c}({\omega_{1}=1})\right)$ (1)
In the section II-A and II-B we discuss the formulation of the state of the driver. Then in section II-C we discuss the driver state aware repositioning recommendation model.
II-A Taxi Driver’s Preference to Repositioning to a Region
Assuming that the driver $c\in\mathcal{C}$ has rejected the system recommendation, i.e., $\omega_{1}=0$, the driver will next reposition to his preferred destination, with the probability of moving to the region $j\in\mathcal{R}$ given as follows:
$\mathbb{P}_{c}(\{\omega_{2}=j\}|\{\omega_{1}=0\})=\mathcal{L}_{cj}$ (2)
where $\mathcal{L}_{cj}\in[0,1]$ is defined as follows:
$\mathcal{L}_{cj}=\frac{\sigma_{c}(\mathbf{z_{j}})}{\sum_{i\in\mathcal{R}}\sigma_{c}(\mathbf{z_{i}})}\hskip 7.11317pt\text{and}\hskip 7.11317pt\sigma_{c}(\mathbf{z_{j}})=\frac{1}{1+e^{-\mathbf{w_{c}}\cdot\mathbf{z_{j}}}}$ (3)
Logistic function $\sigma_{c}$ maps a vector of features [^1] The features that have been used in this study, to learn the driver’s repositioning preferences are listed in Fig. 4 of Section IV. $\mathbf{z_{j}}$ associated with region $j\in\mathcal{R}$ to a score which lies in the interval $[0,1]$. This score is indicative of the the attractiveness of region $j\in\mathcal{R}$ for driver $c\in\mathcal{C}$. Parameter $\mathbf{w_{c}}\in\mathbb{R}^{|\mathbf{z_{j}}|+1}$ is unique to each driver $c\in\mathcal{C}$ and is learned by separately fitting a logistic function to the historic repositioning decisions of each driver using the dataset [23].
Alternately, if the driver accepts the system recommendation, i.e., $\omega_{1}=1$, the driver will move to the recommended destination. Once, the driver has accepted the system recommendation, there is no uncertainty in his selection of the destination region.
$\mathbb{P}_{c}({\{\omega_{2}=j\}}|\{\omega_{1}=1\})=x_{cj}$ (4)
where,
$x_{cj}=\begin{cases}1~{}&{\text{$c\in\mathcal{C}$ is recommended to $j\in\mathcal{R}$}}\\ 0~{}&{\text{$c\in\mathcal{C}$ is not recommended to $j\in\mathcal{R}$}}.\end{cases}$ (5)
Note, $x_{cj}\in\{0,1\}$ is a decision variable, that must be calculated to achieve an optimal dispatch solution. Each taxi driver is recommended only one destination region $j\in\mathcal{R}$:
$\sum_{j\in\mathcal{R}}x_{cj}=1\hskip 2.84544pt\forall c\in\mathcal{C}.$ (6)
II-B Taxi Drivers’ Confidence in the Recommender System
Every driver, $c\in\mathcal{C}$ has an estimate of the probability of success associated with repositioning to region $j\in\mathcal{R}$ when recommended by the system, i.e., $\theta_{r}=\mathbb{P}_{c}(\{\omega_{3}=1\}|\{\omega_{2}=j\},\{\omega_{1}=1\})$, and following his own preference, i.e., $\theta_{p}=\mathbb{P}_{c}(\{\omega_{3}=1\}|\{\omega_{2}=j\},\{\omega_{1}=0\})$. Driver’s estimate of the recommender system’s success probability and following his preference are assumed to be beta distributed and denoted as $\mathbf{\Theta_{r}(c)}\sim\text{Beta}(\alpha_{r}(c),\beta_{r}(c))$ and $\mathbf{\Theta_{p}(c)}\sim\text{Beta}(\alpha_{p}(c),\beta_{p}(c))$ respectively. Beta distribution has a support of $(0,1)$ and the conjugate symmetry property, which makes them suitable for modeling belief distributions [24].
Outcome of the $k^{th}$ recommendation to driver $c\in\mathcal{C}$ is denoted as $y_{c}(k)\in\{0,1\}$ and defined as follows:
$y_{c}(k)=\begin{cases}1,&\text{if driver $c$ gets allocated following $k^{th}$recc.}\\ 0,&\text{if driver $c$ remains idle following $k^{th}$ recc.}\end{cases}$ (7)
After observing the outcomes of $k$ recommendations, the belief distribution of recommender system’s success probability $\mathbf{\Theta_{p}(c)}$ is updated using the Beta distribution’s conjugate updating rule:
$\mathbf{\Theta_{r}(c)}$ $\mid y_{c}(1),\dots,y_{c}(k)\sim$ (8)
$\text{Beta}\left(\alpha_{0}+\epsilon_{1}\sum_{i=1}^{k}y_{c}(i),\hskip 5.69046pt\beta_{0}+\epsilon_{0}\left(k-\sum_{i=1}^{k}y_{c}(i)\right)\right)$
Here, $\epsilon_{0},\epsilon_{1}$ are the weights used for updating the parameters of the confidence model. The belief distribution $\mathbf{\Theta_{p}(c)}\sim\text{Beta}(\alpha_{p}(c),\beta_{p}(c))$ is similarly updated when the driver follows his own preference.
Each successful recommendation increases the driver’s confidence in the system, moving the estimate right, while each failure decreases it, moving the estimate left. Weights, $\epsilon_{0},\epsilon_{1}$ decide how rapidly the curve shifts left and right respectively. Figure 2 shows how a beta distribution evolves with successive successful recommendations
Figure 2: Posterior updates ($\epsilon_{0}=1,\epsilon_{1}=1$) after six successful recommendations. The red circle marks the true success probability ($\theta_{r}=0.8$), and blue dots represent the driver’s estimates $\mathbb{E}[\mathbf{\Theta_{r}(c)}]$.
We model the driver’s decision-making process as being probabilistic. The driver accepts the system’s recommendation with a probability $\mu(c)$, which is calculated using the Thompson sampling [17] based approach where $M$ samples are drawn one-by-one from the distributions $\mathbf{\Theta_{r}(c)}$ and $\mathbf{\Theta_{p}(c)}$ and compared. Let $\hat{\theta}_{r}$ denote the number sample drawn from $\mathbf{\Theta_{r}(c)}$ that were greater than the samples drawn from $\mathbf{\Theta_{p}(c)}$ and $\hat{\theta}_{p}$ being the vice versa. The probability of a driver accepting the rebalancing recommendation is then defined as:
$\mu(c)=\mathbb{P}_{c}(\{\omega_{1}=1\})=\frac{\hat{\theta}_{r}}{\hat{\theta}_{r}+\hat{\theta}_{p}}=\frac{\hat{\theta}_{r}}{M}$ (9)
Usage of this definition reflects exploration with risk aversion and how the driver weighs the risk of failure. If the probability $\hat{\theta}_{r}$ is high, the driver is more confident that the system will succeed, and they are more likely to follow the recommendation. Conversely, if $\hat{\theta}_{r}$ is low, the driver is less confident in the system and less likely to accept the recommendation, yet the driver’s action selection remains non-deterministic thus incorporating the adherence uncertainty.
II-C Driver State Aware Repositioning Recommendation Model
The proposed repositioning recommendation model is used to provide recommendations to the idle-standing taxi drivers. Decision variable $x$ as specified in equation (5) is used to calculate the optimal repositioning recommendations that will maximize the expected profit accumulated by successful allocations of the drivers in the planning horizon.
II-C1 Supply Vector
Supply vector $\mathbf{s}\in\mathbb{Z}^{|\mathcal{R}|}$ is a random vector consisting of the number of taxi drivers that will be available in each region after the repositioning is completed based on the drivers’ choices and element $\mathbf{s_{j}}$ denotes the number of taxi drivers available in the region $j\in\mathcal{R}$ in the planning horizon $\mathcal{H}$.
Random variable $\mathbf{X_{cj}}$ is a defined on the probability space associated with the driver $c\in\mathcal{C}$, which is denoted as $(\Omega,\mathcal{F},\mathbb{P}_{c})$ such that:
$\mathbf{X_{cj}}=\begin{cases}1~{}&{\text{with probability}\hskip 2.84544pt\mathbb{P}_{c}(\{\omega_{2}=j\})}\\ 0~{}&{\text{with probability}\hskip 2.84544pt1-\mathbb{P}_{c}(\{\omega_{2}=j\})}\end{cases}$ (10)
Where, $\mathbb{P}_{c}(\{\omega_{2}=j\})$ is the total probability of the driver $c\in\mathcal{C}$ moving to region $j\in\mathcal{R}$:
$\mathbb{P}_{c}(\{\omega_{2}=j\})$ $=\mu(c)\cdot x_{cj}+(1-\mu(c))\cdot\mathcal{L}_{cj}$ (11)
To determine the total number of taxis that will be available in each region, we need to consider the probabilistic behavior of each taxi driver and aggregate it across all drivers.
$\mathbf{s_{j}}=\sum_{c\in\mathcal{C}}\mathbf{X_{cj}}$ (12)
Since, $\mathbf{X_{cj}}$ are independent Bernoulli random variables, the total taxi count $\mathbf{s_{j}}$ follows a Poisson binomial distribution [25], which is a generalization of the binomial distribution where the probability of success is not the same for each trial. The probability mass function $P(\{s_{j}=b\})$ of the Poisson binomial distribution is given by :
$\sum_{\begin{subarray}{c}C_{b}\subseteq\mathcal{C}\end{subarray}}\bigg{(}\prod_{c\in C_{b}}\mathbb{P}_{c}(\{\omega_{2}=j\})\prod_{c\notin C_{b}}(1-\mathbb{P}_{c}(\{\omega_{2}=j\}))\bigg{)}$ (13)
where $C_{b}$ is a subset of $\{1,2,...,m\}$ with $b$ elements.
Obtaining the probability distribution of the Poisson binomial distribution is computationally hard, due to the necessity of evaluating all possible subsets of drivers. This complexity arises from the fact that the number of subsets grows exponentially with the number of drivers $m$. Specifically, the probability mass function involves summing over $\binom{m}{k}$ terms, making it impractical for large values of $m$. Moreover, unlike the binomial distribution, the Poisson binomial distribution lacks a simple closed-form expression for its probability mass function, further complicating its use in analytical studies and practical applications.
Despite the computational complexity, we can derive the expected value of $\mathbf{s_{j}}$ due to the linearity of expectation and assuming independence of the Bernoulli trials. The expected value of $\mathbf{s_{j}}$ is given by:
$\mathbb{E}[\mathbf{s_{j}}]$ $=\sum_{c=1}^{m}\mathbb{E}[\mathbf{X_{cj}}]=\sum_{c\in C}\mu(c)\cdot x_{cj}+(1-\mu(c))\cdot\mathcal{L}_{cj}$ (14)
II-C2 Demand Vector
Demand vector $\mathbf{d}\in\mathbb{Z}^{|\mathcal{R}|}$ is a random vector consisting of the number of requests that will appear in each region in the planning horizon. Demand is assumed to be a Gaussian random variable, whose parameters are unknown and need to be estimated.
$\mathbf{d_{j}}\sim\mathcal{N}\left(\nu_{j},\sigma_{j}^{2}\right)=\underbrace{\nu_{j}}_{\text{forecast}}+\underbrace{\mathcal{N}(0,\sigma_{j}^{2})}_{\text{error}}$ (15)
First, a function $g_{j}:\mathbf{\tilde{d}_{j}}\to\nu_{j}$ is learned using a gradient-boosted tree-based regression model [26] for each region $j\in\mathcal{R}$ using the historical demand data [23]. This function maps a vector $\mathbf{\tilde{d}_{j}}$, that contains tuples of demand and covariate information of previous planning horizons to the forecast:
$g_{j}(\mathbf{\tilde{d}_{j}})=\nu_{j}$ (16)
where,
$\mathbf{\tilde{d}_{j}}=\bigg{[}\big{(}d_{j}(h-1),\hat{d}(h-1)\big{)},\ldots,\big{(}d_{j}(h-w),\hat{d}(h-w)\big{)}\bigg{]}$
Each variable $d_{j}(h-t)$ denotes the realization of the random variable $\mathbf{d_{j}(h-t)}$, i.e., the number of passenger demands that appeared in the region $j\in\mathcal{R}$ in the planning horizon $h-t$ and $\hat{d}(h-t)$ denotes the covariate information, which contains three additional features, namely i) hour of the day, ii) day of the week, and iii) day of the month. Walk-forward validation [27] is used to refine the trained model as the new data corresponding to the current time step is available.
Parameter $\sigma_{j}^{2}$ of the error distribution $\mathcal{N}(0,\sigma_{j}^{2})$ is estimated using the residuals from the test dataset, and is defined as the expected value of the squared residuals:
$\sigma_{j}^{2}=\mathbb{E}[(\mathbf{d_{j}}-\nu_{j})^{2}]$ (17)
II-C3 Travel Time Matrix
Travel time matrix $\mathbf{T}\in\mathbb{R}^{n\times n}$ is a random matrix, where $\mathbf{T_{ij}}$ is the time it will take to travel from region $i\in\mathcal{R}$ to region $j\in\mathcal{R}$. $\mathbf{T_{ij}}$ is modelled as a Gaussian ransom variable:
$\mathbf{T_{ij}}\sim\mathcal{N}(\tau_{ij},\varepsilon_{ij}^{2})$ (18)
Function $h:\mathbf{\tilde{T}_{ij}}\to\tau_{ij}$ is learned using a gradient-boosted tree-based regression for the inter-region pairs $(i,j)\in\mathcal{R}\times\mathcal{R}$ using the historical trips data [23]. The function maps the feature vector $\mathbf{\tilde{T}_{ij}}$ that contains i) historic travel times between the region $i$ to $j$, ii) distance between the region $i$ to $j$, iii) hour of the day, and iv) day of the week, to the expected value of the travel time:
$h(\mathbf{\tilde{T}_{ij}})=\tau_{ij}$ (19)
We model the error distribution as Gaussian $\mathcal{N}(0,\varepsilon_{ij}^{2})$ and the parameter, $\varepsilon_{ij}^{2})$ is learned using the residuals from the test dataset, and is defined as the difference between the realization of random variable $\mathbf{T_{ij}}$ and the predicted demand value:
$\varepsilon_{ij}^{2}=\mathbb{E}[(\mathbf{T}_{ij}-\tau_{ij})^{2}].$ (20)
Expected travel time for $c\in\mathcal{C}$, standing in region $i\in\mathcal{R}$ for region $j\in\mathcal{R}$ is denoted as:
$\mathcal{T}_{cj}=\tau_{ij}$ (21)
To ensure that the taxis recommended to go to a particular destination $j\in\mathcal{R}$ reach the destination within the planning horizon, the travel time forecasted using the travel time prediction model $\mathcal{T}_{cj}$ for taxi driver $c\in\mathcal{C}$, must be no more than the length of the duration of the planning horizon:
$x_{cj}\cdot\big{(}\mathcal{T}_{cj}-|\mathcal{H}|\big{)}\leq 0$ (22)
Further, to avoid over-saturation, i.e., dispatching more taxis than that are required, we use the approach similar to [7], and define an upper limit of the total number of repositioning recommendations.
$\sum_{c\in\mathcal{C}}x_{cj}\leq\rho\cdot\mathbb{E}[\mathbf{d_{j}}]$ (23)
II-C4 Objective Function
The expected number of taxis that will be allocated in region $j\in\mathcal{R}$ in the planning horizon are given as follows:
$\mathbb{E}\big{[}\min\big{(}\mathbf{s_{j}},\mathbf{d_{j}}\big{)}\big{]}$ (24)
where $\mathbf{s_{j}}$ is a Poisson binomial distributed random variable and $\mathbf{d_{j}}$ is a normally distributed random variable as discussed previously. Obtaining a closed-form expression for $\mathbb{E}\big{[}\min\big{(}\mathbf{s_{j}},\mathbf{d_{j}}\big{)}\big{]}$ is practically intractable due to the computational complexity of deriving the probability distribution of $\mathbf{s_{j}}$. Therefore, we rely on the expected supply $\mathbb{E}[\mathbf{s_{j}}]$, and define an approximation to the expected allocation. This is determined using the expected supply $\mathbb{E}[\mathbf{s_{j}}]$ and expected demand $\mathbb{E}[\mathbf{d_{j}}]$:
$\xi_{j}=\min\big{(}\mathbb{E}[\mathbf{s_{j}}],\mathbb{E}[\mathbf{d_{j}}]\big{)}.$ (25)
The objective is to maximize the expected revenue, i.e., the anticipated earnings from successful allocations. The expected earnings is defined as the product of expected number of allocations $\xi_{j}$ and the expected ride fare $\mathcal{P}_{j}$ (calculated using  [23]) in each region and is given by:
$\sum_{j\in\mathcal{R}}\xi_{j}\cdot\mathcal{P}_{j}$ (26)
The cost $\mathcal{Q}_{cj}$ of moving a taxi from its current location, say region $i\in\mathcal{R}$ to region $j\in\mathcal{R}$, is calculated similarly to  [8], based on the distance between the regions, the average taxi mileage, and the fuel cost. For a driver, the profit is defined as the difference between the earning and costs. Note that the cost of operation is not considered in the objective function, this is because the taxi companies operating in the Uber style model do not bear the transportation cost, rather it the driver who has to bear this cost. The company gets a fixed share of the earnings, and from the companies point of view maximizing overall earnings is a desirable objective.
II-C5 Repositioning Recommendation Model
The repositioning recommendation model (27) generates an optimal taxi-destination match, ensuring the taxis are dispatched to destinations with a higher chance of getting allocated during the planning horizon. Using Eq. 6, Eq. 22, and Eq. 23 as the constraints and expected profit maximization as the objective function, where $\xi_{j}$ is given by Eq. 25 the proposed model can be written as:
Problem II.1.
Optimal Taxi Rebalancing Recommendation (OTRR).
$\max_{x}$ $\sum_{j\in\mathcal{R}}\min\big{(}\mathbb{E}[\mathbf{s_{j}}],\mathbb{E}[\mathbf{d_{j}}]\big{)}\cdot\mathcal{P}_{j}$ (27)
s.t. $\sum_{j\in\mathcal{R}}x_{cj}=1\hskip 2.84544pt\forall c\in\mathcal{C}$
$\sum_{c\in\mathcal{C}}x_{cj}\leq\rho\cdot\mathbb{E}[\mathbf{d_{j}}]\hskip 2.84544pt\forall j\in\mathcal{R}$
$x_{cj}\cdot\big{(}\mathcal{T}_{cj}-|\mathcal{H}|\big{)}\leq 0\hskip 2.84544pt\forall c\in\mathcal{C},j\in\mathcal{R}$
$x_{cj}\in\{0,1\}\hskip 2.84544pt\forall c\in\mathcal{C},j\in\mathcal{R}.$
The proposed model can be converted to a linear program by using an auxiliary variable $\mathcal{Z}_{j}$ such that, the objective function now becomes the product of $\mathcal{Z}_{j}$ and $\mathcal{P}_{j}$ and two new constraints are introduced:
$\mathcal{Z}_{j}\leq\mathbb{E}[d_{j}]\hskip 2.84544pt\forall j\in\mathcal{R}$ (28)
$\mathcal{Z}_{j}\leq\mathbb{E}[s_{j}]\hskip 2.84544pt\forall j\in\mathcal{R}.$
Note that $\mathbb{E}[d_{j}]=\nu_{j}$ is the output of the demand prediction model as defined by Eq. 16, and $\mathbb{E}[s_{j}]$ is a function of driver preference probability $\mathcal{L}_{cj}$ Eq. 3, driver confidence level Eq. 9, and decision variable $x_{cj}$ Eq. 5 as defined by Eq. 14. Additionally, the constraint Eq. 6 can be relaxed with an inequality and decision variable $x_{cj}$ can also be relaxed to $x_{cj}\in[0,1]$. Incorporating the above mentioned relaxations, the model Eq. 27 can be converted to:
Problem II.2.
Linearized OTRR:
$\max_{x,\mathcal{Z}}$ $\sum_{j\in\mathcal{R}}\mathcal{Z}_{j}\cdot\mathcal{P}_{j}$ (29)
s.t. $\sum_{j\in\mathcal{R}}x_{cj}\leq 1\hskip 2.84544pt\forall c\in\mathcal{C}$
$\sum_{c\in\mathcal{C}}x_{cj}\leq\rho\cdot\mathbb{E}[\mathbf{d_{j}}]\hskip 2.84544pt\forall j\in\mathcal{R}$
$x_{cj}\cdot\big{(}\mathcal{T}_{cj}-|\mathcal{H}|\big{)}\leq 0\hskip 2.84544pt\forall c\in\mathcal{C},j\in\mathcal{R}$
$\mathcal{Z}_{j}\leq\nu_{j}\hskip 2.84544pt\forall j\in\mathcal{R}$
$\mathcal{Z}_{j}\leq\sum_{c\in C}\mu(c)\cdot x_{cj}+(1-\mu(c))\cdot\mathcal{L}_{cj}\hskip 2.84544pt\forall j\in\mathcal{R}$
$x_{cj}\in[0,1]\hskip 2.84544pt\forall c\in\mathcal{C},j\in\mathcal{R}$
$\mathcal{Z}_{j}\in\mathbb{R}_{+}\hskip 2.84544pt\forall j\in\mathcal{R}.$
III Algorithm
The proposed algorithm takes the following inputs, i) confidence parameters for each driver: $\alpha_{r}(c),\beta_{r}(c)$, $\alpha_{p}(c),\beta_{p}(c)$, ii) a preference prediction model for each driver: $\sigma_{c}$, iii) features associated with each region for preference prediction: $\mathbf{z_{j}}$, iv) features associated with each region for demand prediction: $\mathbf{\tilde{d_{j}}}$, v) a demand prediction model of each region:$g_{j}$, vi) features associated with each region pair for travel time prediction $\mathbf{\tilde{T}_{ij}}$, vii) a travel time prediction model: h, and viii) expected profits associated with each region $j\in\mathcal{R}$ in planning horizon: $\mathcal{P}_{j}$; and returns the repositioning recommendations $x_{cj}$ for each idle-standing driver. This algorithm is executed every $\mathcal{H}$ mins. at the beginning the planning horizon.
Algorithm 1 Driver State Aware Reposition Recommendation
1:  Input: $\{\alpha_{r}(c),\beta_{r}(c),\alpha_{p}(c),\beta_{p}(c)\forall c\},\{\sigma_{c}\forall c\},\{\mathbf{z_{j}}\forall j\},$
2:  $\{\mathbf{\tilde{d}_{j}}\forall j\},\{g_{j}\forall j\},\{\mathbf{\tilde{T}_{ij}}\forall i,j\},h,\{\mathcal{P}_{j}\forall j\},$
3:  Output: $x$
4:  Evaluate preference of idle drivers using Eq. 3
5:  for $c\in\mathcal{C}$ do
6:     for $j\in\mathcal{R}$ do
7:        $\mathcal{L}_{cj}\leftarrow\text{get\_driver\_preference}(\sigma_{c},\mathbf{z_{j}})$
8:     end for
9:  end for
10:  Evaluate confidence-level of idle drivers using Eq. 9
11:  for $c\in\mathcal{C}$ do
12:     $\mu(c)\leftarrow\text{get\_driver\_confidence}(\alpha_{r}(c),\beta_{r}(c),\alpha_{p}(c),\beta_{p}(c))$
13:  end for
14:  Evaluate expected demand in each region using Eq. 16
15:  for $j\in\mathcal{R}$ do
16:     $\nu_{j}\leftarrow\text{get\_expected\_demand}(g_{j},\mathbf{\tilde{d}})$
17:  end for
18:  Evaluate expected travel time using Eq. 19
19:  for $i\in\mathcal{R}$ do
20:     for $j\in\mathcal{R}$ do
21:        $\tau_{ij}\leftarrow\text{get\_expected\_travel\_time}(h,\mathbf{\tilde{T}_{ij}})$
22:     end for
23:  end for
24:  Evaluate travel time for drivers using Eq. 21
25:  for $c\in\mathcal{C}$ do
26:     $i\leftarrow\text{get\_current\_region\_of\_driver}(c)$
27:     for $j\in\mathcal{R}$ do
28:        $\mathcal{T}_{cj}\leftarrow\text{get\_travel\_time\_for\_driver}(\tau_{ij})$
29:     end for
30:  end for
31:  Evaluate repositioning recommendations by solving (29)
32:  $x\leftarrow\text{get\_recommendations}(\mathcal{L},\mathbf{\mu},\mathbf{\nu},\mathcal{T},\mathcal{P})$
33:  for $c\in\mathcal{C}$ do
34:     if $y_{c}=1$ then
35:        $\alpha_{r}(c)\leftarrow\alpha_{r}(c)+\epsilon_{1}$
36:     else
37:        $\beta_{r}(c)\leftarrow\beta_{r}(c)+\epsilon_{0}$
38:     end if
39:  end for
40:  Evaluate Metric Eq. 30, Eq. 31, Eq. 32, Eq. 33
41:  Return: $x$
IV Case Studies
In this section, we discuss the experiments that we have performed to test the proposed driver state aware repositioning recommendation algorithm using large-scale fleet operations consisting of fleet size ranging from 2000 to 8000 taxis. We explain the operational details of the proposed driver preference and confidence-aware taxi repositioning recommendation model and discuss the components that will become the key ingredients of a realistic test-bed used for simulation case studies. We develop a simulated taxi network based on real data [23] from New York City. [^2] Specifically, we have used three months of data from [23] for training and testing the prediction models. The repositioning recommendation model (27) was triggered every $\mathcal{H}=60$ minutes in simulation. To generate the road network, we leveraged an approach similar to our prior work [8], in which city Manhattan was partitioned into regions denoted by the set $\mathcal{R}$.
Figure 3: This figure shows a taxi and three potential repositioning destinations along with the shortest route, and the boundaries of the regions. An optimal destination is selected by the model and recommendation is given to the driver, who is free to accept or reject the recommendation.
We first determine the centers of each region using the k-means clustering algorithm. Secondly, we leverage the road network to define a directed graph with the nodes representing the road junctions, and the edges representing the links between the road junctions. We then use Dijkstra’s shortest path algorithm to calculate the shortest distance route between the region centers. We denote $\mathcal{\delta}$ to be a matrix that stores the shortest path distance $\mathcal{\delta}_{ij}$ between region $i,j\in\mathcal{R}$. taxi drivers pick up passengers from one region and drop them off at the other region. After dropping off the customer, a driver has a choice to either keep waiting or reposition to a new region. Fig. 3 shows an idle-standing driver and the top 3 potential destinations which will maximize his likelihood of getting the ride in next $\mathcal{H}$ minutes.
IV-A Taxi Driver’s Preference to Reposition
To evaluate the performance of the driver preference prediction problem, we use accuracy as the metric of evaluation which is defined in the context of top-$k$ destinations. We predict the top-$k$ destinations for a driver by selecting the regions with the highest $\mathcal{L}_{cj}$ values. If the driver’s actual choice of destination falls within these $k$ choices, we label the prediction as correct. The fraction of total number of correct predictions out of total number of predictions made for the driver is defined as the accuracy. As shown in right plot in Fig. 4, the accuracy with which the proposed model can pinpoint ($k=1$) the exact choice of driver destination is around 73%. Note that this accuracy varies from driver to driver due to a varied degree of randomness in each driver’s decision-making process and 73% is the average accuracy value for all the drivers. As the value of $k$ increases, the accuracy increases as well, e.g., we can predict that a driver will head to one of the two predicted destinations ($k=2$) with an accuracy of 83%, one of three predicted destinations ($k=3$) with an accuracy of 87% and one of four predicted destinations ($k=4$) with an accuracy of 90%. For the repositioning recommendation system, it is unnecessary to predict the exact destination of a driver, as this is neither possible with absolute certainty due to the inherent random choices of drivers, nor desirable, as we want drivers to remain more exploratory, i.e., they have a non-zero probability of moving to more than one region. This exploratory nature will ensure that the drivers can be steered to destinations in an optimized manner avoiding over supply. What is desired is a probability distribution over the regions, representing a driver’s choice of the repositioning destinations, which can be used to estimate the expected supply distribution (14).
Figure 4: i)The figure on the left shows the feature importance plot highlighting that search distance plays the highest role in predicting taxi driver repositioning preference, ii) The figure on the right shows the accuracy improvement as the number of destination increases.
The dataset reported in [23] was used to extract the features for training the driver preference prediction model and Fig. 4 shows the features used for predicting the driver preference. As a first step, we need to identify the unique drivers in the dataset, which can be done by filtering out the drivers whose “medallion” matches with the “hack license”. A unique driver is the one for whom a trail of pick-ups and drop-offs which is coherent with time can be established. This means that every pickup is later than the previous drop-off. This information has been provided because in dataset [23], multiple drivers can have same “hack license” at the same time. Since, dataset [23] contains only the pick-up and drop-off data, a driver’s repositioning choice is defined as the next region of pickup after a drop-off. Each driver’s preference is learned by training a separate logistic regression model. The drivers whose choice can be predicted with the highest accuracy are selected. These drivers are chosen to be representative drivers, and a group of drivers are assumed to have the same preference as that of the representative driver in the group, e.g., if there are 2000 drivers in the simulation, then groups of 20 drivers may have same preference distribution as their representative driver. This assumption is taken because of practical limitation of data availability, e.g., in dataset [23] there are only 925 unique drivers and not all of them have a predictable preference. Each driver preference model is trained using the features shown in the left plot in Fig. 4. It was found that the top most feature to predict if a driver will move to the destination or not was the distance between the two regions. It is obvious that a driver will not go very far away in search of the passengers. Second influential feature is the median trip distance at the destination. This is because longer trip distance translates to higher profits. Note that the labels or targets in the logistic regression model are {1,0}, where 1 implies that the driver will go the destination. Since, raw data only contains the data points corresponding to the selected destinations (label=1), we augment the training data with label=0 for all the other destinations. The corresponding features such as search distance, median trip distance etc, can be easily extracted from the raw data. So, for each label=1 data point, we have 65 data points each with the label=0, since there are 66 regions in total. Once the logistic regression model is trained, we can evaluate its performance using accuracy evaluation, which measures how often the model correctly predicts {0,1}. However, this is not a good metric due to the significant imbalance between the label=0 and label=1 data points. The abundance of label=0 data points makes it easier to reject a potential destination, leading to misleadingly high accuracy values—around 99% for almost all drivers. So, we choose a modified accuracy as a metric of evaluation as detailed previously and shown in the right figure in Fig. 4 which plots the prediction accuracy vs number of predicted destinations.
Figure 5: i) Plot on top left shows the distribution of the distance that a driver is willing to travel in search of passengers, ii) Plot on top right shows the distribution of the number of pickups that leads to the driver stay in the same region, iii) Plot on the bottom left shows the distribution of the trip distance at the destination iv) Plot on the bottom right shows the distribution of the hours in which the driver chooses to reposition or stay in the same region
The impact of individual factors on driver’s repositioning choice is shown in the Fig. 5. First plot shows the distribution of distance of destination region for two decisions, i.e., to reposition or to wait in the same station. The red plot shows the distribution of the distance that the driver is ready to travel in search of passengers. The blue plot shows the distribution of the distance that the driver will not travel in search of passengers. As seen in the plot, the driver rarely travels more than 5 km in search of passengers. The second plot shows how the distribution of the pickups in a region impacts a driver’s decision to stay in the same region or move to the another region. As seen in the plot the drivers are aware of the upcoming demands in the region they are standing because the driver repositions himself if the demand falls below a certain threshold. The red plot is the distribution of the pickups for which the driver chooses to move to another region. The blue plot shows the distribution of the pickups in the region for which the driver stays in the same region. As seen in the third plot the drivers are aware of the median trip distance at the destination. This can be seen as a spike in the probability distribution (blue plot) in which the drivers don’t move to the regions where median trip distance is almost zero. Finally, w.r.t. hour of the day, although there isn’t a clear preferred hour, there is a slight preference to stay in same region during evenings.
IV-B Taxi Driver’s Confidence on Recommender System
As the driver keeps interacting with the recommender system, the driver’s belief is updated using Eq. 8 in each time step. We create three scenarios based on the values of the weights $\epsilon_{0},\epsilon_{1}$:
  1. 1.
    Neutral $\epsilon_{0}=\epsilon_{1}$ : This means that the driver’s confidence will increase and decrease at the same rate due to successful as well as failed recommendations.
  2. 2.
    Pessimistic $\epsilon_{0}<\epsilon_{1}$ : This means that the driver’s confidence will decrease at a higher rate due to failed recommendations as compared to the rate of increasing in the confidence level due to successful recommendations.
  3. 3.
    Optimistic $\epsilon_{0}>\epsilon_{1}$ : This means that the driver’s confidence will increase at a higher rate due to successful recommendations as compared to the rate of decrease in the confidence level due to failed recommendations.
Fig. 6 shows how the confidence of a driver evolves over time when interacting with the recommender system. The confidence of the drivers in the optimistic case grows the fastest, followed by the neutral drivers, and slowest by the pessimistic drivers. Note that for the pessimistic drivers, the confidence level doesn’t converge to 1, rather it converges to 0.6. This means that there shall be some uncertainty at the steady-state in a driver’s response to recommendations. All the drivers initially start with a low confidence in the recommender system (0.2). In each time step, a driver is given a recommendation to reposition. If a driver is recommended to move to other destination, and the driver gets allocated within same time-step, then the recommendation is considered successful, else the recommendation is considered as failed.
Figure 6: This figure shows the evolution of the average confidence level of the drivers over time for three class of drivers.
IV-C Demand Prediction
The demand prediction problem was formulated as a time-series forecasting problem. Features used for training an XGBoost model [26] are explained in Section II-C2. To determine the optimal look-back window size, i.e., the number of previous hours of demand to consider for training, the window size was defined as a parameter alongside three other model parameters: (i) the number of estimators, (ii) max depth, and (iii) learning rate. Grid search-based approach was employed to identify the best model parameters. As baselines, two pre-trained deep learning-based large language models were used: CHRONOS [28], and MOIRAI [29].
Figure 7: Error Analysis: i) Fig. on the left shows the distribution of the error for three models, ii) Fig. on the right shows the distribution of absolute errors for three models
The performance evaluation of three demand forecasting models was conducted using key error metrics, as summarized in Table I. XGBoost demonstrated the best overall performance, achieving the lowest median absolute error of 3.91, compared to MOIRAI (5.03) and CHRONOS (8.49), indicating superior central tendency accuracy. At the 25th percentile, XGBoost also had the smallest error (0.97) followed by MOIRAI (1.33) and CHRONOS (1.98), showing better performance for relatively easier predictions. For the 75th percentile, XGBoost maintained its lead with an error of 11.02, outperforming MOIRAI (14.05) and CHRONOS (26.43), reflecting robustness for challenging predictions. At the 95th percentile, XGBoost continued to perform best, achieving an error of 31.79 compared to MOIRAI (40.66) and CHRONOS (77.29), highlighting its effectiveness in handling extreme outlier scenarios. Overall, XGBoost emerged as the most reliable model, followed by MOIRAI with moderate performance, while CHRONOS had the highest errors, indicating weaker predictive capability.
Metric XGBoost MOIRAI CHRONOS
Median Absolute Error 3.91 5.03 8.49
25th Percentile Error 0.97 1.33 1.98
75th Percentile Error 11.02 14.05 26.43
95th Percentile Error 31.79 40.66 77.29
TABLE I: Error statistics of demand prediction models.
A spatial error analysis, summarized in Table II, further evaluates the forecasting performance across different regions. The analysis reveals the highest median errors (top 10 regions) ranging from 10.75 to 16.07 across regions, with percentage errors remaining relatively low, between 6.38% and 10.33%.
Region ID Median Error Median Demand %age Error
161 16.07 155.50 10.33
186 15.32 150.50 10.18
236 13.83 183.50 7.54
142 13.69 165.00 8.29
237 13.65 214.00 6.38
162 13.12 152.50 8.60
230 12.75 124.50 10.24
48 11.47 116.00 9.89
170 10.76 133.50 8.06
68 10.75 116.50 9.22
TABLE II: Spatial Error Analysis: Median absolute errors and percentage errors of top 10 regions (out of 66 regions in Manhattan) with region IDs as defined in the dataset [23].
IV-D Travel Time Prediction
The figure on the left in Fig. 8 shows the probability distribution of the error in travel time prediction. The median absolute error in predicting the inter-station travel time is approximately 2 minutes and the 99th percentile of error is approximately 11 minutes. The error in the travel time prediction impacts the reachability of the idle-sanding taxis as defined by Eq. 20. Reachability is defined as the fraction of inter region trips that can be served within the planning horizon (in this case 60 minutes) [5]. Duration of the planning horizon plays a critical role in determining the reachability of the taxi drivers. The figure on the right in Fig. 8 shows the reachability (percentage) vs planning horizon (minutes) plot. The black curve shows that a planning horizon approximately 30 minutes is sufficient to ensure that a taxi is able to reposition between any two regions within the planning horizon. Further, the red curve shows that in addition to repositioning, a taxi driver can also finish a trip if the planning horizon is further increased to 60 minutes. To achieve this, we only consider the trips that are no longer than 30 minutes. Such trips account for 95% of all the trips in the raw data. This helps us defining the time step to be 60 minutes in duration, and the benefits are two fold, first is that the planning horizon duration is sufficiently large so that the demands can be forecasted with high accuracy in terms of error percentage(small planning horizon duration lead to sparse time-series which are difficult to forecast) [30], and second is that it ensures that each taxi can reposition and serve a customer within the planning horizon, thus it makes it possible to discretize the simulation into time-steps, in which the sequence of operations in Algorithm 1 can be executed.
Figure 8: i) Plot on the left shows the probability distribution of the absolute error, ii) Plot on the right shows the reachability %age as the times step duration increase
IV-E Evaluation Metrics
To evaluate the performance of the proposed algorithm we perform a simulation over 7 days. Since each time step is 60 minutes in duration, the total number of time steps in the simulation are 168. For each time step, we record the information related to the performance of the repositioning recommender system and the following metrics are evaluated at the end of the simulation:
IV-E1 Allocation
Allocation is defined as the fraction of drivers that got allocated in a time step. If a driver is allocated in the time $k\in\mathcal{K}$ the reward $y_{c}(k)=\sum_{j\in\mathcal{R}}y_{cj}(k)$ is 1, else if the driver remains idle, then the reward is 0 as defined in (7). Allocation is evaluated using the following expression:
$\frac{1}{|\mathcal{K}|}\frac{1}{|\mathcal{C}|}\sum_{k\in\mathcal{K}}\sum_{c\in\mathcal{C}}y_{c}(k)$ (30)
IV-E2 Driver Profit
Profit for each driver $c\in\mathcal{C}$ is defined as the difference between the earnings made by the driver in time step $k\in\mathcal{K}$ and the rebalancing cost and allocation cost. As defined earlier, $x_{cj}(k)$ is the recommendation decision variable, and the earning of the driver $c\in\mathcal{C}$ be $\mathcal{P}_{cj}(k)$. Also, $y_{cj}(k)$ is the allocation decision variable and $\mathcal{Q}_{cj}(k)$ is the cost of moving the vehicle. We use the average driver profit as a metric which is evaluated as:
$\frac{1}{|\mathcal{K}|}\frac{1}{|\mathcal{C}|}\sum_{k\in\mathcal{K}}\sum_{c\in\mathcal{C}}\sum_{j\in\mathcal{R}}y_{cj}(k)\cdot(\mathcal{P}_{cj}(k)-\mathcal{Q}_{cj}(k))-x_{cj}(k)\cdot\mathcal{Q}_{cj}(k)$ (31)
IV-E3 Driver Confidence
Median confidence of the drivers at the end of the simulation is used as a measure to evaluate the impact of the performance of the repositioning recommender system on a driver’s confidence which is evaluated as follows:
$Median(\mu(\mathcal{|\mathcal{K}|}))$ (32)
IV-E4 Met Demand
Met demand is the fraction of the total number of requests that were served. The served demand is also same the number of allocated taxis and is evaluated as:
$\frac{\sum_{k\in\mathcal{K}}\sum_{i\in\mathcal{R}}\sum_{c\in\mathcal{C}_{i}(k)}\sum_{j\in\mathcal{R}}y_{cj}(k)}{\sum_{k\in\mathcal{K}}\sum_{i\in\mathcal{R}}d_{i}(k)}$ (33)
TABLE III: Model Performance Analysis: Metric Evaluation
Fleet Size Metric Model (Optimistic) Baseline (Optimistic) Model (Neutral) Baseline (Neutral) Model (Pessimistic) Baseline (Pessimistic)
2000 Allocation 0.786 0.783 0.784 0.784 0.772 0.804
Driver Profit 11.242 10.065 11.231 10.094 11.076 10.519
Met Demand 0.374 0.373 0.373 0.373 0.368 0.383
Confidence 1.000 1.000 1.000 1.000 0.900 0.920
4000 Allocation 0.722 0.693 0.723 0.692 0.709 0.661
Driver Profit 9.742 9.020 9.748 9.043 9.640 8.800
Met Demand 0.687 0.659 0.688 0.659 0.675 0.629
Confidence 1.000 1.000 1.000 1.000 0.940 0.470
6000 Allocation 0.641 0.603 0.641 0.600 0.634 0.518
Driver Profit 8.440 7.973 8.453 7.946 8.422 6.951
Met Demand 0.915 0.861 0.915 0.857 0.904 0.740
Confidence 1.000 1.000 1.000 1.000 0.810 0.120
8000 Allocation 0.509 0.496 0.513 0.492 0.507 0.399
Driver Profit 6.725 6.518 6.771 6.486 6.737 5.347
Met Demand 0.969 0.944 0.976 0.936 0.966 0.760
Confidence 1.000 1.000 1.000 1.000 0.360 0.030
Figure 9: This figure shows the performance comparison of the Linearized OTRR (29) with baseline model [7] on metrics defined in IV-E.
IV-F Baseline Model
For comparing the performance of the proposed rebalancing model, we use the model by [7] as a baseline. This agent-level model has proved to be efficient in balancing supply-demand when the supply-agents are assumed to be absolutely adherent. It is agnostic to taxi driver’s preferences as well as impact of recommendations on the driver’s adherence to recommender system. Additionally, this model uses the same set of constraints as (27), so the primary difference is the two model is the objective function. The baseline optimization model is as follows:
Problem IV.1.
Baseline model Waller et al. 2018 [7]
$\max_{x}$ $\sum_{c\in\mathcal{C}}\sum_{j\in\mathcal{R}}x_{cj}\cdot\mathbb{E}[\mathbf{d_{j}}]\cdot(1-\frac{\mathcal{T}_{cj}}{\mathcal{H}})$ (34)
s.t. $\sum_{j\in\mathcal{R}}x_{cj}\leq 1\hskip 2.84544pt\forall c\in\mathcal{C}$
$\sum_{c\in\mathcal{C}}x_{cj}\leq\rho\cdot\mathbb{E}[\mathbf{d_{j}}]\hskip 2.84544pt\forall j\in\mathcal{R}$
$x_{cj}\cdot\big{(}\mathcal{T}_{cj}-|\mathcal{H}|\big{)}\leq 0\hskip 2.84544pt\forall c\in\mathcal{C},j\in\mathcal{R}.$
Both the models are run within the same environment, the only difference being the rebalancing model. The drivers are initialized with 20% confidence level. The simulation in run for a period of 7 days (168 time steps) after which the driver confidence converges as shown in Fig 6. At each time step, we extract the evaluation metrics: Allocation (30), driver profit (31), confidence (32), and the met demand (33). Then the two models are compared based on these metrics.
IV-G Model Performance
This section highlights key observations from the results in Table III, focusing on allocation efficiency, driver profit, met demand, and system confidence under optimistic, neutral, and pessimistic settings. The percentage improvement for each metric is calculated as the average improvement across all fleet sizes and scenarios. The proposed rebalancing model demonstrates superior performance across multiple fleet sizes and evaluation metrics compared to the baseline model under most scenarios as shown in the Fig. 9.
The proposed model consistently achieves higher allocation rates than the baseline across all fleet sizes, particularly in neutral and pessimistic scenarios. For instance, at a fleet size of 4000 under pessimistic conditions, the proposed model achieves an allocation rate of 0.709, outperforming the baseline’s 0.661. Similarly, for a fleet size of 6000, the model achieves 0.634 compared to the baseline’s 0.518. On average, the proposed model improves allocation efficiency by 6.3% compared to the baseline.
Driver profit is significantly higher for the proposed model across all fleet sizes and scenarios. Notably, in the optimistic scenario for a fleet size of 2000, the proposed model yields the hourly driver profit of 11.242 compared to 10.065 for the baseline. This trend continues for larger fleets, with the proposed model maintaining higher profitability for all three classes of the drivers. The profit differences become more pronounced in case of the pessimistic drivers. For instance, at a fleet size of 8000 under pessimistic conditions, the proposed model achieves a profit of 6.737, significantly outperforming the baseline’s 5.347. The overall improvement in driver profit is 11.8% more as compared to the baseline.
The proposed model outperforms the baseline in meeting demand across all fleet sizes, especially under pessimistic scenarios where operational challenges of moving the drivers are most severe. At a fleet size of 6000 under pessimistic conditions, the proposed model meets 90.4% of the demand, compared to 74.0% for the baseline. Even in the optimistic scenario for smaller fleets (e.g., 2000), the proposed model performs slightly better than the baseline with a demand satisfaction rate of 0.374 compared to 0.373. On average, the proposed model improves demand satisfaction by 4.5%.
The confidence metric indicates system reliability and driver trust. While both models converge to a confidence level of 100% in optimistic and neutral scenarios, the proposed model demonstrates resilience under pessimistic conditions. For example, at a fleet size of 4000 under pessimistic conditions, the confidence score for the proposed model remains high at 0.940, significantly outperforming the baseline’s 0.470. For larger fleets like 8000, the baseline’s confidence drops dramatically to 0.030, whereas the proposed model maintains a reasonable level of 0.360. On average, the proposed model improves confidence by 14.7% under pessimistic conditions.
V Conclusions
In the paper we have formulated a driver preference and confidence aware taxi rebalancing model that can provide the repositioning recommendations to the taxi drivers. The proposed model also provides a quantitative way to study the dynamics of the driver’s confidence-level on the recommender system, and how incorporating driver preferences and confidence levels can enhance the performance in terms of the driver profits, and fleet utilization. Extensive experiments show that the proposed model performs better than the model that is agnostic to these attributes of the taxi driver, and hence can be used to provide better repositioning recommendations. For the future work, we plan to incorporate the impact of surge-pricing on the vehicle rebalancing efficiency considering the uncertainty in adherence by the taxi drivers.