### v2

# Vehicle Rebalancing Under Adherence Uncertainty

Abstract
Ride-hailing systems often suffer from spatiotemporal supply-demand imbalances, largely due to the independent and uncoordinated actions of drivers. While existing fleet rebalancing methods offer repositioning recommendations to idle drivers to improve service efficiency, they typically assume full driver compliance: an unrealistic premise in practice. We propose an Adherence-Aware Vehicle Rebalancing (AAVR) framework that explicitly models and addresses uncertainties in driver adherence, stemming from individual behavioral preferences and dynamic trust in the recommender system. Our approach integrates (i) region-specific XGBoost models for demand forecasting, (ii) a network-level XGBoost model for inter-region travel time prediction, (iii) driver-specific logit models to capture repositioning preferences, and (iv) driver-specific Beta-Bernoulli Bandit models with Thompson Sampling to track and update each driver’s confidence in the system over time. These elements are incorporated into a novel optimization framework that generates adherence-aware repositioning recommendations. To enable real-time implementation, we further develop a linearized version of the AAVR model. Extensive simulations on the NYC taxi dataset demonstrate that AAVR significantly outperforms four state-of-the-art adherence-agnostic baselines, achieving a 28% increase in served demand, a 22.7% reduction in customer wait times, a 28.6% increase in platform earnings and a 26% gain in driver profits on average.
Index Terms:
vehicle rebalancing, human factors in cyber-physical systems.
I Introduction
Recent urbanization trends have led to increased travel and its associated externalities, such as increasing levels of congestion worldwide. In this context, the reliance on private cars for personal mobility is becoming increasingly impractical and unsustainable. Ride-hailing services have emerged as an alternative, providing Mobility-on-Demand (MoD) services, but they raise concerns related to the exploitation of public resources, equity, profitability, and, importantly, scalability. Specifically, the travel demand for such services is spatio-temporally asymmetrically distributed (e.g., commuting to the downtown in the morning), making overall operations imbalanced and extremely sensitive to disturbances [1]. Vehicle rebalancing has emerged as a key solution to address these challenges. While several models have been proposed, most assume full compliance from taxi drivers with rebalancing recommendations. In practice, however, drivers may choose to accept or ignore suggestions based on personal preferences or their trust in the system. In this work, we propose an Adherence Aware Vehicle Rebalancing (AAVR) model that explicitly accounts for such behavioral uncertainty. By leveraging historical trip data, AAVR learns driver-specific adherence patterns and incorporates them into the rebalancing optimization process, resulting in more robust and realistic recommendations.
I-A Literature Review
Two main classes of approaches have been proposed in the literature [1]. First, researchers have studied mathematical optimization models [2, 3, 4, 5, 6, 7, 8, 9], where upcoming demand is assumed to be known either exactly or approximated using demand prediction models. The optimization model then calculates rebalancing policies based on predicted demands. Second, researchers have employed reinforcement learning techniques to learn taxi rebalancing policies directly from data [10, 11, 12, 13]. Furthermore, from the decision-making perspective, models can be categorized into two classes, i) aggregate models [2, 3, 4, 5], and ii) agent-level models  [6, 7, 8, 9]. The former generate decisions at the fleet-level, determining how many taxis should be moved from one region to another. However, such models do not consider the state of individual taxi drivers, making it challenging to incorporate driver-level objectives such as fairness and preferences. In contrast, agent-level models possess the desired granularity, and offer greater flexibility and control over individual taxis, allowing for tailored movements based on specific criteria or constraints. However, the enhanced flexibility comes at a computational cost, due to the many variables involved in tracking each vehicle. One of the seminal works in the domain of the taxi rebalancing problem proposes an aggregate-level model for a station based taxi fleet management [2]. The objective is to find the rate at which idle taxis should be dispatched from one station to the other to ensure that the queue length of customers waiting at the station remains bounded. Based on the formulation proposed by the authors, the idle-standing taxis are distributed among the stations proportionally to the number of awaiting customers at each station. The excess taxis (if any) are evenly distributed among the stations at a minimum traveling cost. The primary drawback of this work is that it doesn’t consider the upcoming demands while generating the rebalancing decision, and the repositioning decisions are made solely based on the number of customers waiting in the queue. Subsequent work [3] introduced a more complex model, by accounting for both awaiting customers and expected future customer requests in the decision-making process. Taxis are assigned to stations to guarantee that each station has at least as many taxis as there are awaiting customers. Furthermore, following the allocation for awaiting customers, any remaining taxis are distributed proportionally based on demand forecasts at each station. A major shortcoming of [3] is that a simplistic travel demand forecasting assuming demand arrival as a Poisson process. Secondly, [3] does not account for the uncertainties in the demand forecasts. Refer to [14] for a detailed analysis on travel demand forecasting, and [15] for robust approaches to handle demand uncertainties in the context of the vehicle rebalancing problem. This is critical, because recommendations might be overly optimistic and cause drivers to head to locations with insufficient demand. This may cause distrust among drivers and hence lead to poor adherence to system recommendations in the future. Authors in [4] further enhanced the aggregate-level taxi rebalancing solution by formulating a multi-period stochastic rebalancing problem considering uncertainties in demand forecasts. The objective of this framework is to ensure that the supply-demand ratio at each station is the same as the overall supply-demand ratio at the network level while minimizing the traveling cost. Formulating the taxi rebalancing problem as a multi-period problem has the advantage of accounting for the downstream effects of decisions made in the planning horizon. As pointed out by the authors, however, this may not always be beneficial due to the underlying uncertainties in the forecasts for mobility patterns. Authors in [5] discuss an upper-confidence-bound based approach to determine optimal parameters, such as the rebalancing frequency, and the length of the planning horizon, to ensure the effective implementation of the aggregate-level models sequentially.
All the models discussed so far are aggregate models and are unable to generate taxi-level decisions. [6] is the agent-level version of the aggregate model introduced by authors in [4], and enables decision-making at the individual taxi level. A notable limitation of this work is its failure to account for uncertainties in travel time and their potential impact on taxi availability at the destination. The authors in [7] greatly improved on this aspect and accounted for the impact of the travel time on taxi availability at the destination station. The availability of a taxi at the destination was represented as a fraction of the planning horizon during which the taxi is expected to be present at the destination, determined using pre-calculated inter-station travel times. Although [6, 7] present agent-level models, they do not incorporate any driver-centric objectives, rather the objective still remains to balance network-level supply and demand. In our previous work [8], we improved on this aspect by incorporating fairness in providing rebalancing recommendations and [9, 12] discussed the impact of charging-related constraints on the taxi rebalancing problem. Beyond vehicle rebalancing, dynamic pricing [16] [17], and matching-integrated rebalancing [18], have also been proposed as methods to mitigate imbalances.
Interestingly, research on the impact of human factors on recommender systems is limited. In particular, little attention is given to the analysis of confidence level of taxi drivers and the impact of the rebalancing strategies suggested by the recommender systems on the adherence in the future. Existing taxi rebalancing models are agnostic to taxi driver’s confidence levels as well as their repositioning preferences. Therefore, the rebalancing recommendations given by the model may not be effective in practice because the driver may discard them if they are not confident about the recommender system or if the recommendation is not aligned with their preference. Driver’s confidence levels are impacted by the outcome of the repositioning recommendation: if the recommendation leads to a higher reward, the confidence level increases and vice-versa. Additionally, at any particular time instant, the confidence level impacts the likelihood of the driver accepting the repositioning recommendation from the system.
The evolution of agent’s confidence during interactions with recommender systems has been extensively studied, with growing evidence that human exploratory behavior aligns with Bayesian decision-theoretic models. Building on the foundation of Thompson Sampling [19], researchers have demonstrated significant empirical gains in large-scale recommendation tasks. Controlled experiments further support this view: human behavior in uncertain multi-alternative decision-making tasks is well captured by Thompson Sampling mechanisms [20, 21]. Recent work such as QCARE [22] extends this model by introducing adaptive exploration decay, outperforming utility-based baselines in capturing how humans balance exploration and exploitation. In parallel, collective decision-making has also been interpreted through this lens: Krafft et al.  [23] model group intelligence as a distributed form of Thompson Sampling, where individuals sample actions based on social priors such as popularity, further reinforcing the cognitive plausibility of this mechanism. Despite these advances, Thompson Sampling-based confidence models remain largely unexplored in the context of taxi driver recommender systems. We address this gap by embedding a generalized Beta-Bernoulli Thompson Sampling process into our model, enabling dynamic representation of driver’s evolving confidence in system-generated rebalancing recommendations.
Prior works [24, 25, 26] have also shown that taxi drivers display a preference in terms of their passenger-finding strategies. [26] discussed the driver cruising strategies between the trips and highlights that taxi drivers tend to make reasonable choices between rebalancing and parking, heading to high-demand locations based on the time of day. Authors in [24] used a multinomial logit model to learn the probability distribution over the driver’s choice to head to one of the neighboring zones. Furthermore, authors in [25] presented a model that explores the behavior of taxi drivers when searching for passengers under uncertain conditions and three distinct strategies were studied: random search, maximum anticipated pick-up probability search, and maximum anticipated revenue search. A work that closely relates to our work is by  [27] which discusses a survey based driver preference aware repositioning recommendation model. A notable difference with this work is that it doesn’t consider the impact of the outcome of the repositioning recommendation on the driver’s probability of accepting the recommendation in the future and assumes a static acceptance probability. In our work, we model the driver’s confidence as a dynamic process, which is impacted by repeated interaction with the recommender system.
In this paper, we present a new detailed model for the behavior of taxi drivers, and tools to compute recommendations for such drivers, so that they can reposition their vehicles to increase their returns. Taxi drivers are modeled as agents characterized by a repositioning preference as well as a confidence level. We leverage the dataset reported in [28] to extract the taxi demands as well as driver preferences. A novel taxi rebalancing model has been proposed which provides repositioning recommendations by predicting the upcoming taxi demands, travel times, driver position, repositioning preference, and confidence in the recommender system. Repositioning recommendations are made sequentially and the agent’s confidence level is updated depending on the reward received after reaching the destination. To learn the taxi driver preference, we first identify the features that can describe the repositioning decision of a driver (e.g., the distance of the driver to the recommended destination, time of the day, day of the week, the expected number of requests at the destination, expected revenue from the ride originating at the destination, etc.). Next, we learn the probability of the driver moving to the destination using a logistic model and identify the top-k most likely destinations. When a driver is standing idle after dropping off a passenger, their next location is a function of the repositioning recommendation, their confidence level, and their preference. The confidence level of a taxi driver is modeled using the Thompson sampling-based Beta-Bernoulli Bandits, inspired by [19]. The performance of the confidence- and preference-aware repositioning model is compared with a state-of-the-art the agnostic repositioning model [7].
The key contributions of this work are detailed as follows:
  1. 1.
    We develop a probabilistic framework to model how a taxi driver makes decisions in response to repositioning recommendations. The model captures the sequence of choices a driver makes - whether to accept or reject the recommendation, which region to move to, and whether they are subsequently allocated a passenger. It accounts for differences in individual driver behavior by assigning personalized probabilities to each decision stage.
  2. 2.
    Building on the proposed probabilistic framework, we design a learning-based architecture to estimate the key components of each driver’s behavioral model from data. Specifically, we use (i) region-specific XGBoost models to forecast passenger demand; and (ii) a network-level XGBoost model to estimate inter-region travel times; - which are used to estimate allocation probabilities, (iii) driver-specific logit models to estimate the conditional probabilities of region selection when the driver rejects system recommendation; and (iv) Beta-Bernoulli bandit models with Thompson Sampling to infer and update each driver’s probability of accepting recommendations over time, based on their interaction history.
  3. 3.
    We formulate an Adherence Aware Vehicle Rebalancing (AAVR) model that integrates driver-specific adherence probabilities into a constrained optimization problem. The original model involves a non-linear objective that jointly maximizes expected allocations and minimizes travel time. We derive an equivalent linearized formulation by introducing auxiliary variables, enabling the problem to be solved efficiently as a mixed-integer linear program (MILP).
  4. 4.
    We analyze the proposed strategy in a simulated taxi network derived from real taxi data from a state-of-the-art New York City dataset, and illustrate its effectiveness in comparison to four adherence-agnostic state-of-the-art taxi rebalancing models and show that these models perform sub-optimally as compared to the proposed model in the uncertain adherence setting.
The rest of this manuscript is organized as follows. In Section II a rigorous mathematical formulation of the proposed preference- and confidence-aware rebalancing model is introduced, followed by the detailed presentation of the proposed algorithm for the rebalancing recommender system in Section III. We present detailed simulation-based case studies in Section IV, and draw conclusions in Section V.
II Mathematical Formulation
[…]
Figure 1: This figure illustrates the sequential decision-making process of a taxi driver $c\in\mathcal{C}$, showing the driver’s decision to accept or reject a system recommendation, the subsequent movement to a region, and the allocation outcome.
In this section, we present a detailed mathematical formulation of the proposed taxi drivers’ state aware repositioning recommender system. The system consists of a set of regions denoted as $\mathcal{R}$, and a set of idle taxi drivers denoted as $\mathcal{C}$. Drivers transport the passengers from one region to the other. Idle standing drivers receive a repositioning recommendation which they can either accept or reject. The driver moves to the recommended region if he accepts the system recommendation, or moves to his preferred region if he rejects the system recommendation. Upon reaching, the driver either gets allocated to transport a passenger, or remains idle.
This sequential decision-making process for driver $c\in\mathcal{C}$ is modelled as a random experiment with a probability space $(\Omega,\mathcal{F},\mathbb{P}_{c})$ and is shown in Fig. 1. The sample space of the random experiment is $\Omega=\{0,1\}\times\mathcal{R}\times\{0,1\}$, where an outcome is a tuple $\omega=(\omega_{1},\omega_{2},\omega_{3})\in\Omega$. Here, $\omega_{1}=0$ indicates that the driver rejected the system recommendation, while $\omega_{1}=1$ indicates that the driver accepted the system recommendation. Additionally, $\omega_{2}=j$ implies that the driver moved to region $j\in\mathcal{R}$. Finally, $\omega_{3}=1$ implies that the driver was allocated, whereas $\omega_{3}=0$ indicates that the driver remained idle. The $\sigma$-algebra, of the sample space is denoted as $\mathcal{F}$. For brevity, the following notation is used: $\{\omega_{1}=a\}=\{(\omega_{1},\omega_{2},\omega_{3})\in\Omega\mid\omega_{1}=a\}$. Similarly, this notation can be extended to other combinations of $\omega_{1}$, $\omega_{2}$, and $\omega_{3}$ to represent events involving multiple conditions. The probability measure is defined such that $\mathbb{P}_{c}(\{\omega_{1}=1\})$ represents the probability of the driver accepting the system recommendation, while $\mathbb{P}_{c}(\{\omega_{1}=0\})$ represents the probability of the driver rejecting the system recommendation. Furthermore, $\mathbb{P}_{c}(\{\omega_{2}=j\}\mid\{\omega_{1}=0\})$ denotes the probability of the driver moving to destination $j\in\mathcal{R}$ given that the driver has rejected the system recommendation, and $\mathbb{P}_{c}(\{\omega_{2}=j\}\mid\{\omega_{1}=1\})$ denotes the probability of the driver moving to destination $j\in\mathcal{R}$ given that the driver has accepted the system recommendation. Finally, the probability of the driver getting allocated is defined such that if the driver rejected the system recommendation and moved to region $j\in\mathcal{R}$, the probability that the driver remains idle is $\mathbb{P}_{c}(\{\omega_{3}=0\}\mid\{\omega_{1}=0\},\{\omega_{2}=j\})=1-\theta_{p}$, and the probability that the driver gets allocated is $\mathbb{P}_{c}(\{\omega_{3}=1\}\mid\{\omega_{1}=0\},\{\omega_{2}=j\})=\theta_{p}$. Note that the value of $\theta_{p}$ is unique to the driver $c\in\mathcal{C}$. Similarly, if the driver accepted the system recommendation and moved to region $j\in\mathcal{R}$, the probability that the driver remains idle is $\mathbb{P}_{c}(\{\omega_{3}=0\}\mid\{\omega_{1}=1\},\{\omega_{2}=j\})=1-\theta_{r}$, and the probability that the driver gets allocated is $\mathbb{P}_{c}(\{\omega_{3}=1\}\mid\{\omega_{1}=1\},\{\omega_{2}=j\})=\theta_{r}$.
The state of each taxi driver is defined as tuple containing i) probability of the driver going to region $j\in\mathcal{R}$ by following his own preference, and ii) probability of the driver accepting the reposition recommendation from the system:
$\left(\mathbb{P}_{c}({\omega_{2}=j|\omega_{1}=0}),\mathbb{P}_{c}({\omega_{1}=1})\right).$ (1)
In the section II-A and II-B we discuss the formulation of the state of the driver. Then in section II-C we discuss the driver state aware repositioning recommendation model.
II-A Taxi Driver’s Preference to Repositioning to a Region
Assuming that the driver $c\in\mathcal{C}$ has rejected the system recommendation, i.e., $\omega_{1}=0$, the driver will next reposition to his preferred destination, with the probability of moving to the region $j\in\mathcal{R}$ given as follows:
$\mathbb{P}_{c}(\{\omega_{2}=j\}|\{\omega_{1}=0\})=L_{cj},$ (2)
where $L_{cj}\in[0,1]$ is defined as follows:
$L_{cj}=\frac{\sigma_{c}(\mathbf{z_{j}})}{\sum_{i\in\mathcal{R}}\sigma_{c}(\mathbf{z_{i}})}\hskip 7.11317pt\text{and}\hskip 7.11317pt\sigma_{c}(\mathbf{z_{j}})=\frac{1}{1+e^{-\mathbf{w_{c}}\cdot\mathbf{z_{j}}}}.$ (3)
We use a logit model to learn the repositioning preferences of the taxi drivers. Specifically, a logistic function $\sigma_{c}$ maps a vector of features [^1] The features that have been used in this study, to learn the driver’s repositioning preferences are listed in Fig. 5 of Section IV. $\mathbf{z_{j}}$ associated with region $j\in\mathcal{R}$ to a score which lies in the interval [0,1][0,1][ 0 , 1 ]. This score is indicative of the the attractiveness of region $j\in\mathcal{R}$ for driver $c\in\mathcal{C}$. Parameter $\mathbf{w_{c}}\in\mathbb{R}^{|\mathbf{z_{j}}|+1}$ is unique to each driver $c\in\mathcal{C}$ and is learned by separately fitting a logistic function to the historic repositioning decisions of each driver using the dataset [28].
Alternately, if the driver accepts the system recommendation, i.e., $\omega_{1}=1$, the driver will move to the recommended destination. Once, the driver has accepted the system recommendation, there is no uncertainty in his selection of the destination region:
$\mathbb{P}_{c}({\{\omega_{2}=j\}}|\{\omega_{1}=1\})=\mathbf{x_{cj}},$ (4)
where
$x_{cj}=\begin{cases}1~&{\text{$c\in\mathcal{C}$ is recommended to $j\in\mathcal{R}$}}\\ 0~&{\text{$c\in\mathcal{C}$ is not recommended to $j\in\mathcal{R}$}}.\end{cases}$ (5)
Note, $\mathbf{x_{cj}}\in\{0,1\}$ is a decision variable, that must be calculated to achieve an optimal dispatch solution. Each taxi driver is recommended only one destination region $j\in\mathcal{R}$:
$\sum_{j\in\mathcal{R}}\mathbf{x_{cj}}=1\hskip 2.84544pt\forall c\in\mathcal{C}.$ (6)
II-B Taxi Drivers’ Confidence in the Recommender System
Every driver, $c\in\mathcal{C}$ has an estimate of the probability of success associated with repositioning to region $j\in\mathcal{R}$ when recommended by the system, i.e., $\theta_{r}=\mathbb{P}_{c}(\{\omega_{3}=1\}|\{\omega_{2}=j\},\{\omega_{1}=1\})$, and following his own preference, i.e., $\theta_{p}=\mathbb{P}_{c}(\{\omega_{3}=1\}|\{\omega_{2}=j\},\{\omega_{1}=0\})$. Driver’s estimate of the recommender system’s success probability and following his preference are assumed to be beta distributed and denoted as $\mathbf{\Theta_{r}(c)}\sim\text{Beta}(\alpha_{r}(c),\beta_{r}(c))$ and $\mathbf{\Theta_{p}(c)}\sim\text{Beta}(\alpha_{p}(c),\beta_{p}(c))$ respectively. Beta distribution has a support of (0,1)(0,1)( 0 , 1 ) and the conjugate symmetry property, which makes them suitable for modeling belief distributions [29].
The outcome of the $k^{th}$ recommendation to driver $c\in\mathcal{C}$ is denoted as $y_{c}(k)\in\{0,1\}$ and defined as follows:
$y_{c}(k)=\begin{cases}1,&\text{if driver $c$ gets allocated following $k^{th}$ recc.}\\ 0,&\text{if driver $c$ remains idle following $k^{th}$ recc.}\end{cases}$ (7)
After observing the outcomes of $k$ recommendations, the belief distribution of recommender system’s success probability $\mathbf{\Theta_{p}(c)}$ is updated using the Beta distribution’s conjugate updating rule:
$\mathbf{\Theta_{r}(c)}$ $\mid y_{c}(1),\dots,y_{c}(k)\sim$ (8)
$\text{Beta}\left(\alpha_{0}+\epsilon_{1}\sum_{i=1}^{k}y_{c}(i),\hskip 5.69046pt\beta_{0}+\epsilon_{0}\left(k-\sum_{i=1}^{k}y_{c}(i)\right)\right).$
Here, $\epsilon_{0},\epsilon_{1}$ are the weights used for updating the parameters of the confidence model. The belief distribution $\mathbf{\Theta_{p}(c)}\sim\text{Beta}(\alpha_{p}(c),\beta_{p}(c))$ is similarly updated when the driver follows his own preference.
Each successful recommendation increases the driver’s confidence in the system, moving the estimate right, while each failure decreases it, moving the estimate left. Weights, $\epsilon_{0},\epsilon_{1}$ decide how rapidly the curve shifts left and right respectively. Figure 2 shows how a beta distribution evolves with successive successful recommendations
Figure 2: Posterior updates ($\epsilon_{0}=1,\epsilon_{1}=1$) after six successful recommendations. The red circle marks the true success probability ($\theta_{r}=0.8$), and blue dots represent the driver’s estimates $\mathbb{E}[\mathbf{\Theta_{r}(c)}]$.
We model the driver’s decision-making process as being probabilistic. The driver accepts the system’s recommendation with a probability $\mu(c)$, which is calculated using the Thompson sampling [19] based approach where $M$ samples are drawn one-by-one from the distributions $\mathbf{\Theta_{r}(c)}$ and $\mathbf{\Theta_{p}(c)}$ and compared. Let $\hat{\theta}_{r}$ denote the number sample drawn from $\mathbf{\Theta_{r}(c)}$ that were greater than the samples drawn from $\mathbf{\Theta_{p}(c)}$ and $\hat{\theta}_{p}$ being the vice versa. The probability of a driver accepting the rebalancing recommendation is then defined as:
$\mu(c)=\mathbb{P}_{c}(\{\omega_{1}=1\})=\frac{\hat{\theta}_{r}}{\hat{\theta}_{r}+\hat{\theta}_{p}}=\frac{\hat{\theta}_{r}}{M}.$ (9)
Usage of this definition reflects exploration with risk aversion and how the driver weighs the risk of failure. If the probability $\hat{\theta}_{r}$ is high, the driver is more confident that the system will succeed, and they are more likely to follow the recommendation. Conversely, if $\hat{\theta}_{r}$ is low, the driver is less confident in the system and less likely to accept the recommendation, yet the driver’s action selection remains non-deterministic thus incorporating the adherence uncertainty.
II-C Adherence Aware Vehicle Rebalancing Model
The proposed model generates repositioning recommendations for idle-standing taxi drivers with the goal of maximizing the total number of successful driver-passenger allocations while minimizing the idle cruising time prior to allocation. It explicitly accounts for uncertainty in driver adherence, influenced by both the driver’s confidence in the recommender system and their individual preference for the suggested destination.
II-C1 Supply Vector
Supply vector $\mathbf{s}\in\mathbb{Z}^{|\mathcal{R}|}$ is a random vector consisting of the number of taxi drivers that will be available in each region after the repositioning is completed based on the drivers’ choices and element $\mathbf{s_{j}}$ denotes the number of taxi drivers available in the region $j\in\mathcal{R}$ in the planning horizon $\mathcal{H}$.
Random variable $\mathbf{X_{cj}}$ is a defined on the probability space associated with the driver $c\in\mathcal{C}$, which is denoted as $(\Omega,\mathcal{F},\mathbb{P}_{c})$ such that:
$\mathbf{X_{cj}}=\begin{cases}1~&{\text{with probability}\hskip 2.84544pt\mathbb{P}_{c}(\{\omega_{2}=j\})}\\ 0~&{\text{with probability}\hskip 2.84544pt1-\mathbb{P}_{c}(\{\omega_{2}=j\})}.\end{cases}$ (10)
Where, $\mathbb{P}_{c}(\{\omega_{2}=j\})$ is the total probability of the driver $c\in\mathcal{C}$ moving to region $j\in\mathcal{R}$:
$\mathbb{P}_{c}(\{\omega_{2}=j\})$ $=\mu(c)\cdot\mathbf{x_{cj}}+(1-\mu(c))\cdot L_{cj}.$ (11)
To determine the total number of taxis that will be available in each region, we need to consider the probabilistic behavior of each taxi driver and aggregate it across all drivers.
$\mathbf{s_{j}}=\sum_{c\in\mathcal{C}}\mathbf{X_{cj}}.$ (12)
Since, $\mathbf{X_{cj}}$ are independent Bernoulli random variables, the total taxi count $\mathbf{s_{j}}$ follows a Poisson binomial distribution [30], which is a generalization of the binomial distribution where the probability of success is not the same for each trial. The probability mass function $P(\{s_{j}=b\})$ of the Poisson binomial distribution is given by:
$\sum_{\begin{subarray}{c}C_{b}\subseteq\mathcal{C}\end{subarray}}\bigg{(}\prod_{c\in C_{b}}\mathbb{P}_{c}(\{\omega_{2}=j\})\prod_{c\notin C_{b}}(1-\mathbb{P}_{c}(\{\omega_{2}=j\}))\bigg{)},$ (13)
where $C_{b}$ is a subset of $\{1,2,...,m\}$ with $b$ elements.
Obtaining the probability distribution of the Poisson binomial distribution is computationally hard, due to the necessity of evaluating all possible subsets of drivers. This complexity arises from the fact that the number of subsets grows exponentially with the number of drivers $m$. Specifically, the probability mass function involves summing over $\binom{m}{k}$ terms, making it impractical for large values of $m$. Moreover, unlike the binomial distribution, the Poisson binomial distribution lacks a simple closed-form expression for its probability mass function, further complicating its use in analytical studies and practical applications.
Despite the computational complexity, we can derive the expected value of $\mathbf{s_{j}}$ due to the linearity of expectation and assuming independence of the Bernoulli trials. The expected value of $\mathbf{s_{j}}$ is given by:
$\mathbb{E}[\mathbf{s_{j}}]$ $=\sum_{c=1}^{m}\mathbb{E}[\mathbf{X_{cj}}]=\sum_{c\in C}\mu(c)\cdot\mathbf{x_{cj}}+(1-\mu(c))\cdot L_{cj}.$ (14)
II-C2 Demand Vector
We model the demand vector $\mathbf{d}\in\mathbb{Z}^{|\mathcal{R}|}$ as a random vector representing the number of passenger requests in each region during the planning horizon. Instead of modeling the event-level arrival process (e.g., via a Poisson process), we adopt a forecast-based approach: the expected demand $\nu_{j}$ in each region $j$ is predicted using a supervised regression model (XGBoost), and the uncertainty in this forecast is captured via a Gaussian error model.
Formally, the demand is modeled as:
$\mathbf{d_{j}}\sim\mathcal{N}\left(\nu_{j},\sigma_{j}^{2}\right)=\underbrace{\nu_{j}}_{\text{forecast}}+\underbrace{\mathcal{N}(0,\sigma_{j}^{2})}_{\text{error}}.$ (15)
This formulation reflects a common assumption in forecast-driven optimization, where the prediction error is treated as zero-mean Gaussian noise particularly when probabilistic forecasts are not directly available from the model.
The forecast $\nu_{j}$ is generated by a function $g_{j}:\mathbf{\tilde{d}_{j}}\to\nu_{j}$, trained on historical demand and covariates for each region $j$:
$g_{j}(\mathbf{\tilde{d}_{j}})=\nu_{j},$ (16)
where
$\mathbf{\tilde{d}_{j}}=\bigg{[}\big{(}d_{j}(h-1),\hat{d}(h-1)\big{)},\ldots,\big{(}d_{j}(h-w),\hat{d}(h-w)\big{)}\bigg{]}$
. and $\hat{d}(h-t)$ includes temporal features: hour of the day, day of the week, and day of the month. The variance $\sigma_{j}^{2}$ of the forecast error is estimated using residuals from the test set:
$\sigma_{j}^{2}=\mathbb{E}[(\mathbf{d_{j}}-\nu_{j})^{2}].$ (17)
II-C3 Travel Time Matrix
Travel time matrix $\mathbf{T}\in\mathbb{R}^{n\times n}$ is a random matrix, where $\mathbf{T_{ij}}$ is the time it will take to travel from region $i\in\mathcal{R}$ to region $j\in\mathcal{R}$. $\mathbf{T_{ij}}$ is modelled as a Gaussian random variable:
$\mathbf{T_{ij}}\sim\mathcal{N}(\tau_{ij},\varepsilon_{ij}^{2}).$ (18)
Function $h:\mathbf{\tilde{T}_{ij}}\to\tau_{ij}$ is learned using a gradient-boosted tree-based regression for the inter-region pairs $(i,j)\in\mathcal{R}\times\mathcal{R}$ using the historical trips data [28]. The function maps the feature vector $\mathbf{\tilde{T}_{ij}}$ that contains i) historic travel times between the region $i$ to $j$, ii) distance between the region $i$ to $j$, iii) hour of the day, and iv) day of the week, to the expected value of the travel time:
$h(\mathbf{\tilde{T}_{ij}})=\tau_{ij}.$ (19)
We model the error distribution as Gaussian $\mathcal{N}(0,\varepsilon_{ij}^{2})$ and the parameter, $\varepsilon_{ij}^{2})$ is learned using the residuals from the test dataset, and is defined as the difference between the realization of random variable $\mathbf{T_{ij}}$ and the predicted demand value:
$\varepsilon_{ij}^{2}=\mathbb{E}[(\mathbf{T}_{ij}-\tau_{ij})^{2}].$ (20)
Expected travel time for $c\in\mathcal{C}$, standing in region $i\in\mathcal{R}$ for region $j\in\mathcal{R}$ is denoted as:
$\mathcal{T}_{cj}=\tau_{ij}.$ (21)
To ensure that the taxis recommended to go to a particular destination $j\in\mathcal{R}$ reach the destination within the planning horizon, the travel time forecasted using the travel time prediction model $\mathcal{T}_{cj}$ for taxi driver $c\in\mathcal{C}$, must be no more than the length of the duration of the planning horizon:
$\mathbf{x_{cj}}\cdot\big{(}\mathcal{T}_{cj}-\mathcal{H}\big{)}\leq 0.$ (22)
II-C4 Objective Function
The expected number of taxis that will be allocated in region $j\in\mathcal{R}$ in the planning horizon are given as follows:
$\mathbb{E}\big{[}\min\big{(}\mathbf{s_{j}},\mathbf{d_{j}}\big{)}\big{]}.$ (23)
where $\mathbf{s_{j}}$ is a Poisson binomial random variable and $\mathbf{d_{j}}$ is a normal random variable as discussed previously. Obtaining a closed-form expression for $\mathbb{E}\big{[}\min\big{(}\mathbf{s_{j}},\mathbf{d_{j}}\big{)}\big{]}$ is practically intractable due to the computational complexity of deriving the probability distribution of $\mathbf{s_{j}}$. Therefore, we rely on the expected supply $\mathbb{E}[\mathbf{s_{j}}]$, and define an approximation to the expected allocation. This is determined using the expected supply $\mathbb{E}[\mathbf{s_{j}}]$ and expected demand $\mathbb{E}[\mathbf{d_{j}}]$:
$\xi_{j}=\min\big{(}\mathbb{E}[\mathbf{s_{j}}],\mathbb{E}[\mathbf{d_{j}}]\big{)}.$ (24)
The first of the two objectives in the proposed model is to maximize the total expected allocations across the network. This is desirable from the ride-hailing platform’s perspective because higher expected allocations translate to higher earnings for the platform. We use the following expression to define the total expected allocations:
$J_{1}=\sum_{j\in\mathcal{R}}\min\big{(}\mathbb{E}[\mathbf{s_{j}}],\mathbb{E}[\mathbf{d_{j}}]\big{)}.$ (25)
Additionally, the ride-hailing platforms also care about the passenger waiting time. Although increasing taxi availability which translates to higher allocation also improves the expected waiting time for the passengers, it is possible to achieve the same level of supply distributions by executing different set of recommendations. This requires explicitly ensuring that the drivers closest to the potential pick-up hot spots must be preferred over the drivers that need to travel for longer periods of time. This is achieved by minimizing the total idle cruising time before allocation by the drivers who have been recommended to reposition using the following expression:
$J_{2}=-\sum_{c\in\mathcal{C}}\sum_{j\in\mathcal{R}}\mathbf{x_{cj}}\cdot\mathcal{T}_{cj}.$ (26)
II-C5 Optimization Model
The proposed model generates an optimal taxi-destination match, ensuring the taxis are dispatched to destinations with a higher chance of getting allocated during the planning horizon while ensuring the traversal time to reach the destination is also minimized. Using Eq. 6, Eq. 22, as the constraints and Eq. 25 and Eq. 26 as the objective functions, the proposed Adherence Aware Vehicle Rebalance Model (AAVR) is defined as follows:
Problem II.1.
Adherence Aware Vehicle Rebalancing Model (AAVR).
$\max_{\mathbf{x}\in\{0,1\}^{|\mathcal{C}|\times|\mathcal{R}|}}$ $\sum_{j\in\mathcal{R}}\min\big{(}\mathbb{E}[\mathbf{s_{j}}],\mathbb{E}[\mathbf{d_{j}}]\big{)}-\beta\sum_{j\in\mathcal{R}}\sum_{c\in\mathcal{C}}\mathbf{x_{cj}}\cdot\mathcal{T}_{cj}$ (27)
s.t. $\sum_{j\in\mathcal{R}}\mathbf{x_{cj}}=1\hskip 2.84544pt\forall c\in\mathcal{C}$
$\mathbf{x_{cj}}\cdot\big{(}\mathcal{T}_{cj}-\mathcal{H}\big{)}\leq 0\hskip 2.84544pt\forall c\in\mathcal{C},j\in\mathcal{R}$
The above problem can be converted into a MILP as discussed in Problem IV.1.
III AAVR - Model Architecture and Algorithm
Figure 3: Architecture of the Adherence-Aware Vehicle Rebalancing (AAVR) model. The system comprises four components: (i) an adherence estimator, combining a Beta–Bernoulli bandit with Thompson Sampling and a multinomial logit model, to compute driver-specific acceptance probabilities; (ii) a supply–demand estimator that forecasts adherence-adjusted supply and future demand using region-level XGBoost models; (iii) a travel-time estimator that predicts inter-region travel times via a network-level XGBoost model; and (iv) an optimization module that selects repositioning actions by maximizing expected allocations and minimizing idle travel time.
Figure 3 presents the architecture of the proposed AAVR model, which consists of four core modules: adherence estimation, supply-demand aggregation, travel time prediction, and optimization. The first module estimates the likelihood that each idle-standing driver $\mathrm{DRV}_{i}$ will accept a repositioning recommendation. This adherence probability, is computed by combining two driver-specific models: a Beta-Bernoulli Bandit with Thompson Sampling (BBB-TS), which captures the driver’s confidence in the system’s recommendations based on historical acceptance data (Eq. 9), and a Logit model that estimates the driver’s spatial preferences over regions using regional features (Eq. 3). These outputs are combined to form the adherence probability conditioned on the recommended region (Eq. 11). The second module aggregates supply and demand information across regions. The expected supply in a region is calculated using the number of idle drivers and their adherence probabilities (Eq. 14), while demand is predicted using an XGBoost model trained on historical demand data for each region. The expected number of successful allocations is then derived from these estimates (Eq. 24), and the sum over all regions constitutes the first objective of the optimization problem: maximizing expected total allocations (Eq. 25). The third module estimates inter-region travel times using a network-level XGBoost model trained on historical trip durations. These estimates are used to calculate the time required for each driver to reach any recommended region (Eq. 21), contributing to both the second objective minimizing total idle travel time (Eq. 26) and a feasibility constraint that ensures drivers can reach their assigned regions within the planning horizon (Eq. 22). The final module solves the optimization problem defined in Problem II.1, which balances the trade-off between maximizing expected allocations and minimizing idle travel time while respecting supply, demand, and reachability constraints.
The AAVR - Algorithm 1 is executed at the beginning of each planning horizon, every $\mathcal{H}$ minutes, and takes as input: (i) BBB-TS (Beta Bernoulli Bandit with Thompson Sampling) model for each driver $c\in\mathcal{C}$; (ii) Logit model $\sigma_{c}$ for each driver $c\in\mathcal{C}$; (iii) feature vector $\mathbf{z}_{j}$ for preference prediction and $\mathbf{\tilde{d}}_{j}$ for demand forecasting for each region $j\in\mathcal{R}$; (iv) a demand prediction model $g_{j}$ for each region $j\in\mathcal{R}$; (v) pairwise region features $\mathbf{\tilde{T}}_{ij}$ and a travel time prediction model $h$. Using these inputs, the algorithm produces binary repositioning decisions $\mathbf{x_{cj}}$, indicating whether driver $c$ should reposition to region $j$, with the goal of improving system-level efficiency while accounting for individual driver behavior and spatial preferences. Upon execution, performance metrics such as the total allocation, average wait time, platform profit, and driver profits are calculated to evaluate recommendations.
The overall computational complexity of the AAVR algorithm arises from five main components. First, inter-region travel time estimation incurs a cost of $\mathcal{O}(|\mathcal{R}|^{2})$ due to pairwise computations between all region pairs. Second, demand forecasting contributes $\mathcal{O}(|\mathcal{R}|)$ operations using region-level models. Third, preference evaluation and driver-specific travel time computation require $\mathcal{O}(|\mathcal{C}|\cdot|\mathcal{R}|)$ operations to compute compatibility scores and travel estimates for each driver-region pair. Fourth, the repositioning optimization is formulated as a mixed-integer linear program (MILP) with $|\mathcal{C}|\cdot|\mathcal{R}|$ binary decision variables. While MILPs are NP-hard in the worst case, modern solvers can handle instances of practical size efficiently using branch-and-bound-based algorithms. Finally, updating the confidence model parameters incurs an additional $\mathcal{O}(|\mathcal{C}|)$ cost.
Algorithm 1 Adherence-Aware Vehicle Rebalancing
[…]
2:Output: Rebalancing decision matrix $\mathbf{x}\in\{0,1\}^{|\mathcal{C}|\times|\mathcal{R}|}$
3:Preference & Confidence Estimation
4:for $c\in\mathcal{C}$ do
[…]
6:  for $j\in\mathcal{R}$ do
7:   $L_{cj}\leftarrow\text{get\_driver\_preference}(\sigma_{c},\mathbf{z}_{j})$
8:  end for
9:end for
10:Demand & Travel Time Estimation
11:for $j\in\mathcal{R}$ do
12:  $\nu_{j}\leftarrow\text{get\_expected\_demand}(g_{j},\mathbf{\tilde{d}}_{j})$
13:end for
14:for $c\in\mathcal{C}$ do
15:  $i\leftarrow\text{get\_current\_region}(c)$
16:  for $j\in\mathcal{R}$ do
17:   $\tau_{ij}\leftarrow\text{get\_expected\_travel\_time}(h,\mathbf{\tilde{T}}_{ij})$
18:   $\mathcal{T}_{cj}\leftarrow\tau_{ij}$
19:  end for
20:end for
21:Solve Optimization Problem (Eq. 28)
[…]
23:Bandit Update
24:for $c\in\mathcal{C}$ do
25:  if $y_{c}=1$ then
26:   $\alpha_{r}(c)\leftarrow\alpha_{r}(c)+\epsilon_{1}$
27:  else
28:   $\beta_{r}(c)\leftarrow\beta_{r}(c)+\epsilon_{0}$
29:  end if
30:end for
31:Return: $\mathbf{x}$
IV Case Studies
In this section, we discuss the experiments that we have performed to test the proposed driver state aware repositioning recommendation algorithm using large-scale fleet operations. We explain the operational details of the proposed driver preference and confidence-aware taxi repositioning recommendation model and discuss the components that will become the key ingredients of a realistic test-bed used for simulation case studies. We develop a simulated taxi network based on real data [28] from New York City for the period of January to July 2010.222Specifically, we have used six months of data from [28] for training and one month for testing the prediction models. To generate the road network, we leveraged an approach similar to our prior work [8], in which Manhattan city was partitioned into regions denoted by the set $\mathcal{R}$.
Figure 4: This figure shows a taxi and three potential repositioning destinations along with the shortest route, and the boundaries of the regions. An optimal destination is selected by the model and recommendation is given to the driver, who is free to accept or reject the recommendation.
We first determine the centers of each region using the k-means clustering algorithm. Secondly, we leverage the road network to define a directed graph with the nodes representing the road junctions, and the edges representing the links between the road junctions. We then use Dijkstra’s shortest path algorithm to calculate the shortest distance route between the region centers. We denote $\mathcal{\delta}$ to be a matrix that stores the shortest path distance $\mathcal{\mathcal{D}}_{ij}$ between region $i,j\in\mathcal{R}$. taxi drivers pick up passengers from one region and drop them off at the other region. After dropping off the customer, a driver has a choice to either keep waiting or reposition to a new region. Fig. 4 shows an idle-standing driver and the top 3 potential destinations which will maximize his likelihood of getting the ride in next $\mathcal{H}$ minutes. The repositioning recommendation model (27) was triggered every $\mathcal{H}=5$ minutes to ensure rapid repositioning recommendation and avoid the scenarios where the taxi driver has to wait for the recommendation.
IV-A Taxi Driver’s Preference to Reposition
[…]
Figure 5: i)The figure on the left shows the feature weight plot highlighting that search distance plays the highest role in predicting taxi driver repositioning preference, ii) The figure on the right shows the accuracy improvement as the number of destination increases.
Figure 6: i) Plot on top left shows the distribution of the distance that a driver is willing to travel in search of passengers, ii) Plot on top right shows the distribution of the number of pickups (in 30 mins) that leads to the driver stay in the same region, iii) Plot on the bottom left shows the distribution of the trip distance at the destination iv) Plot on the bottom right shows the distribution of the hours in which the driver chooses to reposition or stay in the same region
The dataset  [28] was used to extract the features for training the driver preference prediction model and Fig. 5 shows the features used for predicting the driver preference. As a first step, we need to identify the unique drivers in the dataset, which can be done by filtering out the drivers whose “medallion” matches with the “hack license”. A unique driver is the one for whom a trail of pick-ups and drop-offs which is coherent with time can be established. This means that every pickup is later than the previous drop-off. This information has been provided because in dataset [28], multiple drivers can have same “hack license” at the same time. Since, dataset [28] contains only the pick-up and drop-off data, a driver’s repositioning choice is defined as the next region of pickup after a drop-off. Each driver’s preference is learned by training a separate logistic regression model. The drivers whose choice can be predicted with the highest accuracy are selected. These drivers are chosen to be representative drivers, and a group of drivers are assumed to have the same preference as that of the representative driver in the group, e.g., if there are 2000 drivers in the simulation, then groups of 20 drivers may have same preference distribution as their representative driver. This assumption is taken because of practical limitation of data availability, e.g., in dataset [28] there are only 925 unique drivers and not all of them have a predictable preference. Each driver preference model is trained using the features shown in the left plot in Fig. 5. It was found that the top most feature to predict if a driver will move to the destination or not was the distance between the two regions. It is obvious that a driver will not go very far away in search of the passengers. Second influential feature is the median trip distance at the destination. This is because longer trip distance translates to higher profits. Note that the labels or targets in the logistic regression model are {1,0}, where 1 implies that the driver will go the destination. Since, raw data only contains the data points corresponding to the selected destinations (label=1), we augment the training data with label=0 for all the other destinations. The corresponding features such as search distance, median trip distance etc, can be easily extracted from the raw data. So, for each label=1 data point, we have 65 data points each with the label=0, since there are 66 regions in total. Once the logistic regression model is trained, we can evaluate its performance using accuracy evaluation, which measures how often the model correctly predicts {0,1}. However, this is not a good metric due to the significant imbalance between the label=0 and label=1 data points. The abundance of label=0 data points makes it easier to reject a potential destination, leading to misleadingly high accuracy values—around 99% for almost all drivers. So, we choose a modified accuracy as a metric of evaluation as detailed previously and shown in the right figure in Fig. 5 which plots the prediction accuracy vs number of predicted destinations.
The impact of individual factors on driver’s repositioning choice is shown in the Fig. 6. First plot shows the distribution of distance of destination region for two decisions, i.e., to reposition or to wait in the same station. The red plot shows the distribution of the distance that the driver is ready to travel in search of passengers. The blue plot shows the distribution of the distance that the driver will not travel in search of passengers. As seen in the plot, the driver rarely travels more than 5 km in search of passengers. The second plot shows how the distribution of the pickups in a region impacts a driver’s decision to stay in the same region or move to the another region. As seen in the plot the drivers are aware of the upcoming demands in the region they are standing because the driver repositions himself if the demand falls below a certain threshold. The red plot is the distribution of the pickups for which the driver chooses to move to another region. The blue plot shows the distribution of the pickups in the region for which the driver stays in the same region. As seen in the third plot the drivers are aware of the median trip distance at the destination. This can be seen as a spike in the probability distribution (blue plot) in which the drivers don’t move to the regions where median trip distance is almost zero. Finally, w.r.t. hour of the day, although there isn’t a clear preferred hour, there is a slight preference to stay in same region during evenings.
IV-B Demand Prediction
The demand prediction problem was formulated as a time-series forecasting problem. Features used for training an XGBoost model [31] are explained in Section II-C2. To determine the optimal look-back window size, i.e., the number of previous hours of demand to consider for training, the window size was defined as a parameter alongside three other model parameters: (i) the number of estimators, (ii) max depth, and (iii) learning rate. Grid search-based approach was employed to identify the best model parameters. As baselines, two pre-trained deep learning-based large language models were used: CHRONOS [32], and MOIRAI [33].
Figure 7: Error Analysis: i) Fig. on the left shows the distribution of the error for three models, ii) Fig. on the right shows the distribution of absolute errors for three models
The performance evaluation of three demand forecasting models was conducted using key error metrics, as summarized in Table I. XGBoost demonstrated the best overall performance, achieving the lowest mean absolute error of 4.05, compared to MOIRAI (6.47) and CHRONOS (29.99), indicating superior average accuracy. It also had the lowest root mean squared error (6.26), suggesting better stability in error distribution compared to MOIRAI (13.06) and CHRONOS (106.96). At the 25th percentile, XGBoost again led with an error of 0.90, followed by MOIRAI (1.32) and CHRONOS (3.81), showing its effectiveness in easier prediction scenarios. For the 75th percentile, XGBoost achieved an error of 5.54, outperforming MOIRAI (8.58) and CHRONOS (32.00), demonstrating robustness under moderately difficult conditions. At the 95th percentile, which highlights performance in extreme cases, XGBoost maintained its lead with an error of 13.23, ahead of MOIRAI (20.74) and CHRONOS (77.00). Overall, XGBoost emerged as the most reliable and consistent model, followed by MOIRAI with moderate performance, while CHRONOS exhibited the highest errors, indicating limited predictive capability.
Metric XGBoost MOIRAI CHRONOS
Mean Absolute Error 4.05 6.47 29.99
Root Mean Squared Error 6.26 13.06 106.96
25th Percentile Error 0.90 1.32 3.81
75th Percentile Error 5.54 8.58 32.00
95th Percentile Error 13.23 20.74 77.00
TABLE I: Error statistics of demand prediction models.
Region ID Median Error Median Demand %age Error
230 6.12 62.0 9.87
161 5.93 59.0 10.05
170 5.69 61.0 9.32
236 5.59 46.0 12.14
237 5.56 49.0 11.35
234 5.37 51.0 10.53
186 5.35 53.0 10.09
162 5.29 51.0 10.38
48 5.23 43.0 12.17
68 4.97 41.0 12.13
TABLE II: Spatial Error Analysis: Median absolute errors and percentage errors of top 10 regions (out of 66 regions in Manhattan) with region IDs as defined in the dataset [28].
A spatial error analysis, summarized in Table II, further evaluates forecasting accuracy across different regions. The analysis highlights the top 10 regions with the highest median errors, which range from 4.97 to 6.12, with percentage errors between 9.32% and 12.17%, indicating a relatively stable performance across the most challenging spatial zones.
IV-C Travel Time Prediction
The figure on the left in Fig. 8 shows the probability distribution of the error in travel time prediction. The median absolute error in predicting the inter-station travel time is approximately 2 minutes and the 99th percentile of error is approximately 11 minutes. The error in the travel time prediction as defined by Eq. 20 impacts the reachability of the idle-sanding taxis. Reachability is defined as the fraction of inter region combinations that can be traversed within the planning horizon [5]. Duration of the planning horizon plays a critical role in determining the reachability of the taxi drivers. The figure on the right in Fig. 8 shows the reachability (percentage) vs planning horizon (minutes) plot. The black curve shows that a planning horizon approximately 30 minutes is sufficient to ensure that a taxi is able to reposition between any two regions of Manhattan within the planning horizon. Further, the red curve shows that in addition to repositioning, a taxi driver can also finish a trip if the planning horizon is further increased to 60 minutes.
Figure 8: i) Plot on the left shows the probability distribution of the absolute error, ii) Plot on the right shows the reachability %age as the times step duration increase
IV-D Analysis of Adherence Aware Vehicle Rebalancing Problem
This section introduces the Linearized AAVR model and benchmarks it against four baseline, state-of-the-art rebalancing strategies. These baselines are implemented to enable a fair and structured comparison across agent-level and aggregate-level rebalancing formulations. The AAVR model uniquely incorporates driver-specific adherence behavior, while the baselines assume full compliance. Specifically, we compare our model (Problem IV.1) against four other ones: B1 [7], B2 as a single-period variant of [6], B3 [3], and B4 as the single-step variant of [18]. Models B1 and B2 are agent-level, producing driver-specific recommmendations directly. In contrast, B3 and B4 operate at the aggregate level and require a post-processing step (Problem IV.6) to generate driver-level actions.
We define the binary variable $\mathbf{x_{cj}}\in\{0,1\}$ as the decision variable for the recommendation at the driver level, where $\mathbf{x_{cj}=1}$ implies that driver $c\in\mathcal{C}$ is recommended to go to station $j\in\mathcal{R}$. Further, $\mathbf{X_{ij}}\in\mathbb{N}$ is defined as the total number of taxis recommended to go from station $i\in\mathcal{R}$ to $j\in\mathcal{R}$ and $\mathbf{Y_{ij}}\in\mathbb{N}$ defines the number of passengers in region $i\in\mathcal{R},$ that are allocated to the drivers in region $j\in\mathcal{R}$. Additionally, let $\mathbf{S}_{i}\in\mathbb{N}$ denote the total number of taxis present in region $i\in\mathcal{R}$ after the execution of rebalancing decisions. Let $\mathbf{Z}_{i}\in\mathbb{N}$ represent the number of taxis that successfully received allocations in region $i$, and $\mathbf{T}_{i}\in\mathbb{N}$ denote the number of unallocated passengers remaining in region $i$ following the rebalancing process. Finally, $\mathcal{D}_{cj}$ denotes the shortest path distance while $\mathcal{T}_{cj}$ denotes the shortest path travel time between driver $c$ and region $j$. The proposed model and the baseline optimization models B1-B4 are defined as follows:
Problem IV.1.
AAVR (Linearized):
$\max_{\mathbf{x},\mathbf{Z}}$ $\sum_{j\in\mathcal{R}}\mathbf{Z_{j}}-\beta\sum_{c\in\mathcal{C}}\sum_{j\in\mathcal{R}}\mathbf{x_{cj}}\cdot\mathcal{T}_{cj}$ (28)
s.t. $\sum_{j\in\mathcal{R}}\mathbf{x_{cj}}=1\hskip 2.84544pt\forall c\in\mathcal{C}$
$\mathbf{x_{cj}}\cdot\big{(}\mathcal{T}_{cj}-\mathcal{H}\big{)}\leq 0\hskip 2.84544pt\forall c\in\mathcal{C},j\in\mathcal{R}$
$\mathbf{Z_{j}}\leq\sum_{c}\mu(c)\cdot\mathbf{x_{cj}}+(1-\mu(c))\cdot L_{cj}\hskip 2.84544pt\forall j\in\mathcal{R}$
$\mathbf{Z_{j}}\leq\nu_{j}\hskip 2.84544pt\forall j\in\mathcal{R}$
The above model is the linearized equivalent of Problem II.1. A new decision variable $\mathbf{Z_{j}}\in\mathbb{N}$ is introduced to replace $\min\left(\mathbf{\mathbb{E}[\mathbf{d_{j}}],\mathbb{E}[\mathbf{s_{j}}]}\right)$ with bounds defined by the newly added constraints, where $\mathbb{E}[\mathbf{s_{j}}]=\sum_{c}\mu(c)\cdot\mathbf{x_{cj}}+(1-\mu(c))\cdot L_{cj}$ and $\mathbb{E}[\mathbf{d_{j}}]=\mathbf{\nu_{j}}$.
Problem IV.2.
Baseline Model 1 (B1)
$\max_{\mathbf{x}}$ $\sum_{c\in\mathcal{C}}\sum_{j\in\mathcal{R}}\mathbf{x_{cj}}\cdot\mathbf{\nu_{j}}\cdot\left(1-\frac{\mathcal{T}_{cj}}{\mathcal{H}}\right)$ (29)
s.t. $\sum_{j\in\mathcal{R}}\mathbf{x_{cj}}\leq 1\hskip 14.22636pt\forall c\in\mathcal{C}$
$\sum_{c\in\mathcal{C}}\mathbf{x_{cj}}\cdot\left(1-\frac{\mathcal{T}_{cj}}{\mathcal{H}}\right)\leq\rho\cdot\mathbf{\nu_{j}}\hskip 14.22636pt\forall j\in\mathcal{R}$
$\mathbf{x_{cj}}\cdot\left(\mathcal{T}_{cj}-\mathcal{H}\right)\leq 0\hskip 14.22636pt\forall c\in\mathcal{C},j\in\mathcal{R}$
The modeling assumptions of the above defined baseline [7] match exactly the proposed model, making it directly implementable. This model uses a parameter $\rho$ to control over-saturation of a region with taxis by bounding the total number of taxis recommended to a region. The authors in [7] do not discuss how to set the value of this parameter. We set $\rho=1$ to use the predicted demand as the target. The model’s objective function is designed to favor regions with higher demands.
Problem IV.3.
Baseline Model 2 (B2)
$\min_{\mathbf{x}}$ $\left\|\frac{\sum_{c}\mathbf{x_{cj}}}{N}-\frac{\mathbf{\nu_{j}}}{\sum_{j}\mathbf{\nu_{j}}}\right\|_{1}-\beta\sum_{c\in\mathcal{C}}\sum_{j\in\mathcal{R}}\mathbf{x_{cj}}\cdot\mathcal{D}_{cj}$ (30)
s.t. $\sum_{j\in\mathcal{R}}\mathbf{x_{cj}}=1\hskip 2.84544pt\forall c\in\mathcal{C}$
$\mathbf{x_{cj}}\cdot\left(\mathcal{T}_{cj}-\mathcal{H}\right)\leq 0\hskip 14.22636pt\forall c\in\mathcal{C},j\in\mathcal{R}$
The above model is a single-step finite horizon variant of the original one [6], which does not consider the travel time constraint. To avoid the scenario that a driver is recommended a destination which is not reachable within the planning horizon, a travel time constraint was included in IV.3.
Problem IV.4.
Baseline Model 3 (B3)
$\min_{\mathbf{X}}$ $\sum_{i}\sum_{j}\tau_{ij}\cdot\mathbf{X_{ij}}$ (31)
s.t. $V_{i}+\sum_{j}\mathbf{X_{ij}}-\sum_{j}\mathbf{X_{ij}}\geq\frac{\mathbf{\nu_{i}}}{\sum_{j}\mathbf{\nu_{j}}}\cdot N\hskip 14.22636pt\forall i\in\mathcal{R}$
$\sum_{j}\mathbf{X_{ij}}\leq V_{i}\hskip 14.22636pt\forall i\in\mathcal{R}$
The above model represents the feedback and proportional predictive rebalancing strategy [3]. The term on the right-hand side of the first constraint is the desired vehicle distribution. Note that the original model does not explicitly include the second constraint, but we include it in this work to ensure that the total number of taxis recommended to reposition in other regions of a region $i\in\mathcal{R}$ must be no more than the taxis that stand in that region. This model has a limitation: it does not consider the travel time feasibility in the case of finite planning horizon. As a result, not all the taxis dispatched to a region reach within the planning horizon. We could not find a direct way to add a constraint to limit such flows because adding a constraint leads to infeasibility due to the first constraint; because it is not always possible to achieve a proportional distribution under restricted flows.
Problem IV.5.
Baseline Model 4 (B4)
$\min_{\mathbf{X},\mathbf{Y},\mathbf{T},\mathbf{S}}$ $\sum_{i}\sum_{j}\mathcal{D}_{ij}\cdot\mathbf{X_{ij}}+\beta\sum_{i}\sum_{j}\mathcal{D}_{ji}\cdot\mathbf{Y_{ij}}+\gamma\sum_{i}\mathbf{T_{i}}$ (32)
s.t. $\sum_{j}\mathbf{Y_{ji}}\leq\mathbf{S_{i}},\sum_{j}\mathbf{Y_{ij}}\leq\mathbf{\nu_{i}},\sum_{j}\mathbf{X_{ij}}\leq V_{i}\hskip 8.5359pt\forall i\in\mathcal{R}$
$\mathbf{T_{i}}=\mathbf{\nu_{i}}-\sum_{j}Y_{ij}\hskip 14.22636pt\forall i\in\mathcal{R}$
$\mathbf{S_{i}}=V_{i}+\sum_{j}\mathbf{X_{ij}}-\sum_{j}\mathbf{X_{ij}}\hskip 14.22636pt\forall i\in\mathcal{R}$
$a_{ij}\cdot\mathbf{X_{ij}}=0,\hskip 14.22636ptb_{ij}\cdot\mathbf{Y_{ij}}=0\hskip 14.22636pt\forall i\in\mathcal{R},j\in\mathcal{R}$
The above model is a single-step variant of [18]. The original model is a multi-step planning problem, but we use its single step variant to ensure uniform modeling assumptions. Unlike [3], this model also has the travel time feasibility embedded in the last constraints:
$a_{ij}=\begin{cases}0~&{\mathcal{T}_{ij}\leq\mathcal{H}}\\ 1~&{\mathcal{T}_{ij}>\mathcal{H}}\end{cases}$ (33)
Similarly, for the allocation decision variable. A taxi in region $j\in\mathcal{R}$ can only be allocated to a passenger in region $i\in\mathcal{R}$ if the travel time is below a certain threshold $w$.
$b_{ij}=\begin{cases}0~&{\mathcal{T}_{ij}\leq w}\\ 1~&{\mathcal{T}_{ij}>w}\end{cases}$ (34)
For this study which uses a node-based approach, we only allow the drivers to allocate a driver to a passenger if the driver and passenger are in the same region (at same node). So $w$ is set to zero because we assume $\mathcal{T}_{ii}=0$, so only $\mathbf{Y_{ii}}$ is allowed to be nonzero.
Problem IV.6.
Model used for conversion of aggregate-level decisions of IV.4 and IV.5 to driver-level decisions.
$\max_{\mathbf{x}}$ $C$ (35)
s.t. $\sum_{j}\mathbf{x_{cj}}\leq 1\hskip 14.22636pt\forall c\in\mathcal{C},j\in\mathcal{R}$
$\sum_{c\in\mathcal{C}_{i}}\mathbf{x_{cj}}=\mathbf{X^{*}_{ij}}\hskip 14.22636pt\forall i\in\mathcal{R}$
Given the optimal rebalancing flows $\mathbf{X^{*}_{ij}}$ from IV.4 or IV.5, driver-level decisions are calculated using the above model, where the objective is any constant value.
Next, we discuss the case studies that we have performed to analyze the performance of the proposed AAVR model. In the first three case studies, we study different variants of two-station problem to elaborate the difference in recommendations by different models illustrating why adherence modeling is critical to ensure an effective vehicle rebalancing model in the context of taxi systems. Next, we illustrate the network level performance of the proposed model and benchmarks.
IV-D1 Case Study 1: Two-Station Problem - Impact of Driver Confidence
In the first case study, we first discuss the underlying logic of the rebalancing decisions generated by the vehicle rebalancing models in a setting where the drivers have a choice to accept or reject the system recommendation. Specifically, we use a two-station setting, say station A and B. Let there be 1000 taxis waiting at station A and 100 pickup requests at station B. The rebalancing system is fully aware of the supply and demand distribution. Further, each driver has a 50% confidence in system, i.e, $\mu(c)=0.5$. If a taxi driver is not recommended to go to station B or recommended to stay at station A, they stay at station A, i.e., $L_{cA}=1.0,L_{cB}=0.0$. If a driver accepts to move to station B, he is assumed to be able to reach it within the planning horizon. The rebalancing model has to decide how many and which taxi drivers should be recommended to go to station B. Logically, the recommender system must recommend 200 drivers to go to station B, so that in expectation 100 drivers move to station B; that would lead to all the demand being served without any waiting drivers. However, the baseline models are not designed considering the compliance uncertainties. Indeed, it is obvious to also note that since none of the baseline models consider the driver confidence, they must recommend the same number of drivers irrespective of the confidence levels of each driver assuming everything else remains the same.
The figure in the upper left of Fig. 9 shows the number of taxi drivers recommended to move to station B. Model B4 recommends 100 drivers to go to station B which in a setting where the confidence-level of all the drivers are completely adherent would be the optimal response if the objective is to serve all the passengers while ensuring that no driver remains idle. However, since the confidence level of each driver is 50%, only 50 out of 100 drivers accept the recommendation as shown in the top right figure in Fig. 9. This leads to 50 passenger being served with no driver remaining idle at station B. In contrast, B2 and B3 recommend all the 1000 drivers to move to station B. This is because both of these models try to match the supply proportion to the demand one. Since the demand proportion between stations A and B is 0: 1, that is, all passenger requests are at station B, all 1000 drivers are recommended to move to station B. In both of the cases B2, B3, all the demand is served by excessive number of drivers remain idle at the station B. The authors in B2 [6] do not discuss how to set the optimal value of the parameter $\beta$ to control the supply in such uncertain compliance setting. B1 uses a parameter $\rho$ to control the over-supply of the region. Additionally, they use a heuristic to not dispatch the exact number of taxis as the demand but rather a higher value depending on the travel time between the two stations. This leads to 500 recommendations when the travel time between station A and B is set to 4 mins and planning horizon is 5 mins. This again leads to fully satisfying the demand but 150 drivers remain idle. Without this heuristic, the model recommends 100 drivers $\rho=1$. So, the model has a flexibility to increase or decrease the flow. However, the authors in [7] also do not discuss how optimal $\rho$ can be selected in such uncertain compliance setting.
Finally, the AAVR model recommends 200 drivers to move to station B, out of which on average 100 drivers accept the recommendation, thus leading to all the demand being fulfilled and no driver remains idle at station B.
Remark 1.
The vehicle rebalancing models that are designed to dispatch the taxis by matching them with the demand such as B4, lead to sub-optimal demand fulfillment especially when confidence among drivers is low. In contrast, models such as B2 and B3 that do not control over-supply of the drivers lead to idle standing drivers, thus causing distrust among the drivers regarding the recommender system. A confidence-aware model such as the AAVR model is more robust to such compliance uncertainties.
Figure 9: Case Study 1: Two-Station Problem - Impact of Driver Confidence on Served Demand and Unallocated Drivers.
IV-D2 Case Study 2: Two-Station Problem - Impact of Driver Preference
In this case study, we investigate the impact of being aware of a taxi driver’s preference to reposition, i.e., the probability of moving to station B if he doesn’t receive any recommendation from the system, or rejects the system recommendation to stay in Station A. In this study, the drivers have a 50% confidence in the recommender system. Unlike case study 1, if the driver did not receive a recommendation or rejects the system’s recommendation to stay at station A, there is a 0.5 probability of the driver moving to station B, i.e., $L_{cA}=0.5,L_{cB}=0.5$.
All baseline models generate the same recommendations because there is no awareness of the driver’s confidence and preference. The situation among the drivers is further aggravated because in addition to the recommended drivers, some drivers reposition to station B on their own, thus leading to more competition for allocation on arriving at station B. The proposed model that is aware of the driver preference interestingly generates no recommendation, anticipating sufficient drivers repositioning on their own. Note that in this case, even if some drivers remain unallocated yet this does not impact the driver’s confidence on the recommender system because the repositioning decision was driver’s own.
Remark 2.
Information about driver preference is critical in designing an effective adherence-based rebalancing model. In cases such as above, the information about the driver preference may lead to fewer recommendations while achieving the objectives of higher served demands without risking the ineffective recommendations.
Figure 10: Case Study 2: Two-Station Problem - Impact of Driver Preference on Served Demand and Unallocated Drivers.
IV-D3 Case Study 3: Two-Station Problem - Driver Selection
In this case study, instead of each driver having a 0.5 confidence in the system’s recommendation, we initialize the driver confidence uniformly between 0 to 1. The driver preference is the same as in case study 1. As seen in Fig. 11 the baseline models as expected generate the same rebalancing decision, However compared to case study 1, in this case the proposed model recommends less than 200 drivers. This is further elaborated by the Fig. 12 which shows that the proposed model unlike the baseline models that are driver confidence agnostic, chooses the drivers with higher confidence in the recommender system while maintaining close to 100 driver in expectation. In this case, the proposed model, and baselines B1, B2, B3 are able to serve all the demands. However, no driver remains unallocated at station B for the proposed model, unlike that for baseline models B1, B2, and B3.
Figure 11: Case Study 3: Two-Station Problem - Impact of random confidence level on driver selection.
Figure 12: Case Study 3: Two-Station Problem - Distribution of the confidence level of the selected drivers.
Remark 3.
Although selecting drivers with higher confidence in the recommender system is desirable from the perspective of the platform, since it increases the likelihood of compliance with the rebalancing recommendations, this preference may introduce a bias that is not always ideal. It is important to note that this represents only a single moment in time. In subsequent case studies, we demonstrate that the proposed model not only accelerates the growth of driver confidence through repeated recommendations, but also improves both platform-centric and passenger-centric performance metrics.
TABLE III: Percentage Improvement Over Baselines B1–B4 Across Different Scenarios
Metric Optimistic Neutral Pessimistic
B1 B2 B3 B4 Avg B1 B2 B3 B4 Avg B1 B2 B3 B4 Avg
Served Demand 9.68 29.26 42.10 28.64 27.42 9.21 31.42 38.69 38.10 29.36 5.00 20.79 37.09 50.78 28.42
Waiting Time -11.49 -24.98 -26.25 -36.83 -24.89 -7.16 -21.79 -22.34 -31.41 -20.68 -13.14 -19.44 -22.98 -34.81 -22.59
Platform Earnings 9.42 27.78 42.62 29.97 27.45 8.99 30.26 38.79 39.75 29.45 5.07 19.77 37.51 53.55 28.97
Driver Profit 6.89 24.93 41.09 25.37 24.57 7.24 28.50 37.80 36.33 27.47 3.05 17.64 36.78 49.55 26.76
Figure 13: Comparison of proposed model with other baselines under optimistic, neutral, and pessimistic scenarios across four performance metrics.
IV-D4 Case Study 4: Network Level Problem
In this case study we do a complete network-level testing of the proposed model in comparison with the baseline models using Manhattan taxi network. The vehicle rebalancing model is executed after every 5 mins, considering the state of the drivers as well forecasted demand in the planning horizon. It is assumed that the demand arrives at the beginning of the time period and stays till the end of time period and drops out at the end of the time period if it remains un-served. First batch of the allocation happens at the beginning of the time period after which the remaining taxis are rebalanced to the nearby stations that can be reached within the planning horizon. Upon arrival the demands at the destination can be served by the taxis on the first come first serve basis. When a taxi is recommended to some station, the driver may accept or reject the recommendation based on the confidence level of the driver. We use 3000 taxis in simulation similar to baseline B1 [7] which leads to almost complete demand fulfillment in absolute adherence setting. All the taxi drivers start with a 50% confidence in the recommender system, suggesting unbiased opinion. The confidence evolves after each recommendation using the Beta-Bernoulli Thompson Sampling based dynamical system as explained in the Driver Confidence section II-B. To evaluate the performance of the proposed model as compared to the baseline models we use four metrics: i) Average Served Demand: This is defined as the total number of passenger requests served on average for a time period of 5 mins. ii) Average Waiting Time: This is defined as the time the passengers had to wait on average for allocation. The max wait time is 5 minutes beyond which the passenger leaves the system, iii) Average Platform Earning: This is defined as the total earnings by the platform. We assume 20% of each ride-share goes to the platform. iv) Average Driver Profit: This is defined as the difference between the ride fare after deduction by the platform and travel cost.
Table III reports the percentage improvement of the Proposed Model over baselines B1–B4 under three demand scenarios: Optimistic, Neutral, and Pessimistic. For Served Demand, the average improvements are 27.42%, 29.36%, and 28.42% respectively across the three scenarios. In terms of Waiting Time, the model achieves average reductions of 24.89%, 20.68%, and 22.59%. For Platform Earnings, the improvements are 27.45%, 29.45%, and 28.97%, while Driver Profit increases by 24.57%, 27.47%, and 26.76%. The gains remain consistently high across all metrics and scenarios, with the largest margins often observed in the pessimistic case. Among the baselines, B1 [7] emerges as the next-best performer across most metrics and scenarios. The empirical evidence shows that the heuristic used by [7] to bias the recommendations towards peak demand areas and recommending more vehicle than the predicted demand works well in this setting with adherence uncertainty, even though the model is not explicitly designed for it.
The results across all scenarios and fleet sizes indicate that the proposed model consistently improves Served Demand, Platform Earnings, and Driver Profit, while also reducing Waiting Time. In the Neutral scenario, Served Demand increased from 28.09% to 29.83% as fleet size grew from 2000 to 4000, while Waiting Time showed a significant reduction from 10.21-10.21- 10.21% to 47.07-47.07- 47.07%. Platform Earnings and Driver Profit also showed stable improvements, peaking at 29.94% and 27.47%, respectively. In the Optimistic scenario, the trend was similar, with Served Demand improving from 26.48% to 29.40%, and Waiting Time reducing sharply from 10.50-10.50- 10.50% to 54.63-54.63- 54.63%. Platform Earnings and Driver Profit reached up to 29.37% and 25.31%, respectively. The Pessimistic scenario yielded the most notable gains, particularly in Served Demand and Platform Earnings, which reached 37.73% and 38.33% at the 4000 fleet size. Despite the challenging conditions, Waiting Time was still significantly reduced from 8.05-8.05- 8.05% to 37.14-37.14- 37.14%. Overall, the results highlight the robustness of the model in improving key performance metrics across different demand scenarios and fleet sizes.
V Conclusions
In the paper we have formulated a driver preference and confidence aware taxi rebalancing model that can provide the repositioning recommendations to the taxi drivers. The proposed model also provides a quantitative way to study the dynamics of the driver’s confidence-level on the recommender system, and how incorporating driver preferences and confidence levels can enhance the performance in terms of the driver profits, and fleet utilization. Extensive experiments show that the proposed model performs better than the model that is agnostic to these attributes of the taxi driver, and hence can be used to provide better repositioning recommendations. For the future work, we plan to incorporate the impact of surge-pricing on the vehicle rebalancing efficiency considering the uncertainty in adherence by the taxi drivers.